### v1

# Watermarking Algorithm for Large Language Models

Abstract
The rapid proliferation of large language models (LLMs) has created an urgent need for reliable methods to detect whether a text is generated by such models. In this paper, we propose SimMark, a posthoc watermarking algorithm that makes LLMs’ outputs traceable without requiring access to the model’s internal logits, enabling compatibility with a wide range of LLMs, including API-only models. By leveraging the similarity of semantic sentence embeddings and rejection sampling to impose detectable statistical patterns imperceptible to humans, and employing a soft counting mechanism, SimMark achieves robustness against paraphrasing attacks. Experimental results demonstrate that SimMark sets a new benchmark for robust watermarking of LLM-generated content, surpassing prior sentence-level watermarking techniques in robustness, sampling efficiency, and applicability across diverse domains, all while preserving the text quality.***The source code of our algorithm can be found at https://github.com/DabiriAghdam/SimMark.
Large Language Models,Watermarking LLMs,Posthoc Watermarking,Sentence-Level Watermarking,Similarity-Based Watermarking, Machine Learning, ICML
1 Introduction
Figure 1: A high-level overview of SimMark detection algorithm. The input text is divided into individual sentences $X_{1}$ to $X_{N}$, which are embedded using a semantic embedding model, with optional PCA for dimensionality reduction. The similarity between consecutive sentence embeddings is computed. Sentences with similarities within a predefined interval $[a,b]$ are considered valid, while those outside are invalid. A soft-$z$-test is performed using the soft count of invalid/partially valid sentences to determine whether the text is watermarked.
The advent of deep generative AIs has made it increasingly important to detect whether a given text is generated by a large language model (LLM). Advanced LLMs, such as GPT-3.5 (OpenAI, 2022), GPT-4 (OpenAI et al., 2023), and Claude 3 (Anthropic, 2024), can produce human-like text at scale and low cost, enabling transformative applications across industries. However, these capabilities also present significant risks, including but not limited to misuse for academic plagiarism, disinformation campaigns and propaganda aimed at manipulating public opinion. For example, the use of AI-generated content in news articles has raised concerns about transparency, accountability, and the potential spread of misinformation (Futurism, 2023). Furthermore, the ability to reliably identify LLM-generated content is essential for enforcing copyright protections and ensuring accountability for such content (Weidinger et al., 2021).
Detecting LLM-generated text poses a unique challenge. These models are explicitly trained to emulate human writing styles, often rendering their outputs indistinguishable from human-authored text. As demonstrated by Kumarage et al. (2023) and Sadasivan et al. (2023), reliably differentiating between human-written and machine-generated text remains an open problem.
One promising approach to addressing this, is the use of imperceptible statistical signatures, or watermarks, embedded within a text. Watermarking imperceptibly alters text such that it remains natural to human readers but enables subsequent detection of its origin (Atallah et al., 2001). Effective watermarking methods must balance the preservation of LLM output quality with robustness against adversarial paraphrasing, where attackers modify the text to evade detection (Krishna et al., 2024). Additionally, watermarks must be resistant to spoofing attacks, wherein adversaries craft non-machine-generated text (often malicious) to falsely trigger watermark detectors (Sadasivan et al., 2023).
In this paper, we introduce SimMark, a robust posthoc sentence-level watermarking algorithm for LLMs based on sentence embedding similarity. The term posthoc here refers to the fact that the watermarking process operates on text after it has been generated by the LLM, rather than modifying the text generation process itself. This approach makes SimMark compatible with a wide range of models, including open-weight LLMs and closed-source proprietary models accessible only via APIs, as it does not require fine-tuning or access to the model’s internal logits. Access to logits is often restricted by API providers due to their potential use in distilling LLMs and leaking proprietary information (Finlayson et al., 2024). SimMark leverages embedding representations from text embedding models to capture semantic relationships between sentences and imposes detectable statistical patterns on sentence similarity through rejection sampling. Specifically, rejection sampling involves querying the LLM multiple times until the similarity between embeddings of consecutive sentences falls within a certain predefined interval. At detection time, these patterns are analyzed using one-proportion soft-$z$-test, a statistical test, to differentiate between human-written and LLM-generated text, as illustrated in Figure 1.
In summary, our contributions are as follows:
  • We introduce a novel watermarking algorithm that achieves state-of-the-art detection performance while maintaining low false positive rates for human-written text.
  • Our approach demonstrates robustness against paraphrasing attacks through semantic-level watermarking and a soft counting mechanism for statistical testing.
  • Compared to existing watermarking methods, SimMark provides a more practical solution that operates without access to LLM logits or fine-tuning, offering efficient and high-quality watermark injection and detection.
The remainder of this paper is organized as follows. Section 2 reviews the background and related work on watermarking techniques for LLMs. Section 3 outlines our methodology. Section 4 describes our experimental setup and presents comparative results, while Section 5 concludes the paper. In the end, we also discuss the potential ethical considerations and societal impacts.
2 Background
In this section, we provide an overview of the foundational concepts related to text generation with LLMs and discuss related works on watermarking techniques for LLMs.
2.1 Autoregressive Decoding of LLMs
An LLM operates over a vocabulary $V$, a set of words or subwords termed as tokens. Let $f:V\rightarrow V$ be an LLM that takes a sequence of tokens $T_{i}=\{t_{1},t_{2},\ldots,t_{i}\}$ as input and generates the next token $t_{i+1}$ as its output.
To generate $t_{i+1}$, the LLM samples it from the conditional probability distribution $P(t_{i+1}|T_{i})$ over the vocabulary $V$. After generating $t_{i+1}$, the updated sequence $T_{i+1}=\{t_{1},t_{2},\ldots,t_{i},t_{i+1}\}$ is fed back into the model $f$, and the process is repeated iteratively to generate the subsequent token $t_{i+2}$. This continues until an end-of-sequence token is generated or the generation reaches a predefined token limit. This sequential process of generating one token at a time based on the previously generated tokens is known as autoregressive decoding.
2.2 Token-Level Watermarking
Token-level watermarking methods embed a statistical signal in the text by manipulating the token generation process (Kirchenbauer et al., 2023; Zhao et al., 2023; Fu et al., 2024). These methods typically alter the probability distribution over $V$ during decoding, subtly biasing the selection of certain tokens to form detectable patterns.
For instance, the KGW algorithm introduced by Kirchenbauer, Geiping, Wen, Katz, Miers, and Goldstein (2023), groups $V$ into green and red subsets pseudo-randomly seeded on the previous token before generating each new token. A predefined positive constant $\delta$ is added to the logits of each token in the green list, increasing their probability of being selected during the sampling step.
At detection, a one-proportion $z$-test is applied to the number of tokens from the green list in the text to determine whether the text contains a watermark. This test compares the observed proportion of green tokens to the expected proportion under the null hypothesis of no watermark, providing a statistical measure to detect even subtle biases introduced by the watermark (Kirchenbauer et al., 2023).
Zhao et al. (2023)’s UNIGRAM-WATERMARK builds upon KGW by fixing the red and green lists instead of pseudo-randomly selecting them, proving that compared to KGW their method is more robust to paraphrasing and editing (Zhao et al., 2023). However, as outlined by Hou et al. (2024), this algorithm can be reverse-engineered, rendering it impractical for high-stakes, real-world applications.
The Semantic Invariant Robust (SIR) watermark in Liu et al. (2023) is also similar to KGW but is designed to be less sensitive to attacks involving synonym replacement or advanced paraphrasing. SIR achieves this by altering the LLM logits based on the semantics of previously generated tokens, using an embedding model to compute semantic representations and training a model that adjusts LLM’s logits based on the semantic embeddings of prior tokens (Liu et al., 2023).
Detection of such watermarks involves analyzing token sequences for statistical signatures that deviate from typical human-generated text. However, token-level watermarks can still be vulnerable to paraphrasing attacks, as rephrasing may disrupt the green and red token lists without altering the overall meaning (Krishna et al., 2024). Moreover, since these methods modify the logits, they directly impact the conditional probability distribution over $V$, potentially affecting the quality of the generated text (Fu et al., 2024).
2.3 Sentence-Level Watermarking
One approach to mitigate the previously mentioned problems is to inject the watermark signal at the sentence-level semantics, making it less vulnerable to paraphrasing attacks. Consider a similar notation for sentence generation using model $f$ (i.e., the LLM) with autoregressive decoding that takes a sequence of sentences $M_{i}=\{X_{1},X_{2},\ldots,X_{i}\}$ and generates the next sentence $X_{i+1}$. The updated sequence of sentences $M_{i+1}=M_{i}\cup\{X_{i+1}\}$ is then used to generate subsequent sentences iteratively.
SemStamp (Hou et al., 2023) employs Locality-Sensitive Hashing (LSH) (Indyk & Motwani, 1998) to pseudo-randomly partition the semantic space of the embedding model into a set of valid and blocked regions, analogous to the green and red subsets in KGW. During rejection sampling, if the representation of a newly generated sentence lies within the valid regions (determined based on the LSH signature of the previous sentence), the sentence is accepted. Otherwise, a new sentence is generated until a valid sentence is produced or the retry limit is reached. Similar to KGW, a one-proportion $z$-test is applied to the number of valid sentences (analogous to green tokens) to determine whether the text contains a watermark.
To improve robustness against paraphrasing, SemStamp used a contrastive learning approach (Hadsell et al., 2006), fine-tuning a semantic embedding model such that the embeddings of paraphrased sentences remain as close as possible to the original sentences. This was achieved by minimizing the distance between paraphrased and original embeddings while ensuring unrelated sentences remained distinct. Additionally, they introduce a margin constraint in the rejection sampling process to reject sentences whose representations lie near the region boundaries. Their results demonstrate that this sentence-level approach improves robustness against paraphrasing.
$k$-SemStamp (Hou et al., 2024) builds upon SemStamp and aims to enhance robustness by partitioning the semantic space using $k$-means clustering (Lloyd, 1982) instead of random partitioning. They claim that in this way, sentences with similar semantics are more likely to fall within the same partition, unlike random partitioning, which may place semantically similar sentences into different partitions, reducing robustness (Hou et al., 2024). However, $k$-SemStamp assumes that the LLM generates text within a specific domain to apply $k$-means clustering effectively (Hou et al., 2024), limiting its applicability in real-world, open-domain scenarios. The generation and detection procedures of $k$-SemStamp remain largely similar to the original SemStamp. In contrast to token-level algorithms, these sentence-level methods do not alter the logits of the LLM, therefore it is expected for their resulting text to be of higher quality (Hou et al., 2023).
Our work, like SemStamp and $k$-SemStamp, is a sentence-level watermarking algorithm; however, it embeds its watermark signature in the semantic similarity of consecutive sentences. It achieves greater generalizability across domains by leveraging an off-the-shelf, general-purpose sentence embedding model without fine-tuning. At the same time, it outperforms state-of-the-art token-level and sentence-level watermarking methods in robustness against paraphrasing while preserving text quality.
3 SimMark: A Similarity-Based Watermark
In this section, we present our proposed framework for watermarking LLMs, detailing both the process of generating watermarked text and its subsequent detection.
Figure 2: Overview of SimMark algorithm. Top: Generation. For each newly generated sentence ($X_{i+1}$), its embedding ($e_{i+1}$) is computed using the Instructor-Large model, optionally applying PCA for dimensionality reduction. The cosine similarity (or Euclidean distance) between $e_{i+1}$ and the embedding of the previous sentence ($e_{i}$), denoted as $s_{i+1}$, is calculated. If $s_{i+1}$ lies within the predefined interval $[a,b]$, the sentence is marked valid and accepted. Otherwise, rejection sampling generates a new candidate sentence until validity is achieved or the iteration limit is reached. Once a sentence is accepted, the process repeats for subsequent sentences. Bottom: Detection (+ Paraphrase attack). Paraphrased versions of watermarked sentences are generated ($Y_{i}$), and their embeddings ($e^{\prime}_{i}$) are computed. The similarity (or distance) between consecutive sentences in the paraphrased text is evaluated. If paraphrasing causes the similarity ($s^{\prime}_{i+1}$) to fall outside $[a,b]$, it is mismarked as invalid. A soft counting mechanism (via function $c(s_{i+1})$ instead of a regular counting with a step function in the interval $[a,b]$) quantifies partial validity based on proximity to the interval bounds, enabling detection of watermarked text via the soft-$z$-test even under paraphrase attacks. It should be emphasized that soft counting is always applied during detection, regardless of whether paraphrasing is present or not, as we cannot assume prior knowledge of paraphrasing.
3.1 Watermarked Text Generation
Similar to SemStamp and $k$-SemStamp (Hou et al., 2023, 2024), SimMark utilizes the embedding representations of the sentences. To compute the embeddings, in contrast with Hou et al. (2023, 2024) that fine-tuned their embedder model (which could make it biased toward a specific paraphrasing model or domain), we employ Instructor-Large (Su et al., 2023), a general-purpose instruction-tuned embedding model. The use of a general-purpose embedding model enables our approach to be more easily adaptable to different domains, and since the generated embeddings are conditioned on a specific instruction, even if a third party knows what embedder model we used, it is still not trivial to reverse-engineer our algorithm without knowing the exact instruction used during generation (the user can also change this instruction frequently), resulting in more robustness against spoofing attacks.
Algorithm 1 Generation Pseudo-Code for SimMark
0:  LLM, embedding model, predefined interval $[a,b]$, maximum iterations $N_{\mathrm{max}}$, optional PCA model
1:  for each generated sentence $i$ ($X_{i}$do
2:     Compute embedding $e_{i}$ for $X_{i}$ using the embedding model.
3:     $n\leftarrow 0$
4:     do
5:         Generate sentence $i+1$ ($X_{i+1}$) using the LLM.
6:         Compute embedding $e_{i+1}$ for $X_{i+1}$ using the embedding model.
7:         Optional: Reduce the dimension of $e_{i}$ and $e_{i+1}$ using the PCA model.
8:         Compute similarity $s_{i+1}\leftarrow sim(e_{i},e_{i+1})$:
$\small s_{i+1}\leftarrow\begin{cases}\tfrac{e_{i}\cdot e_{i+1}}{\|e_{i}\|_{2}\cdot\|e_{i+1}\|_{2}}&\text{ if cosine similarity},\\ \|e_{i}-e_{i+1}\|_{2}&\text{ if Euclidean distance,}\end{cases}$
9:         $n\leftarrow n+1$
10:     while $s_{i+1}{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\notin[a,b]}$ and $n<N_{\mathrm{max}}$
11:     Accept $X_{i+1}$ as valid and continue generating the next sentence. {Here we either reached the $N_{\mathrm{max}}$ or $s_{i+1}{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\in[a,b]}$}
12:  end for
Using Instructor-Large, we compute embeddings for a sentence by passing both the sentence and an instruction (e.g., “Represent the sentence for cosine similarity:” or “Represent the sentence for Euclidean distance:”) to the Instructor model. Then, we calculate the cosine similarity (or Euclidean distance) between the embedding of sentence $i+1$ (denoted as $e_{i+1}$ in Algorithm 1) and the embedding of sentence $i$ (denoted as $e_{i}$). If the computed value lies within the user-defined predefined interval, sentence $i+1$ is considered valid (analogous to the green subset in KGW). Otherwise, we prompt the LLM to generate a new sentence and repeat this procedure until a valid sentence is found or the maximum number of iterations is reached (in this case we accept the last generated sentence, and continue to generating sentence $i+2$), as shown in Algorithm 1. Optionally, we may apply Principal Component Analysis (PCA) method to the embeddings to reduce their dimensionality before calculating the similarity or distance (this is the instruction in this case: “Represent the sentence for PCA:”). The reason for applying PCA is provided in Subsection 3.2. An overview of SimMark generation algorithm is depicted in the top part of Figure 2.
The predefined interval is a hyperparameter chosen a priori by the user based on the distribution of similarities (or distances) between consecutive sentences’ embeddings generated by an unwatermarked LLM and human-written text. The choice of this hyperparameter is critical for the performance of SimMark.
First, if the interval’s width is too small or its position is far from the mean of the similarity distribution, generating sentences within this interval can be challenging or even infeasible for the LLM. Conversely, if the interval’s width is large and centered around the mean of the similarity distribution, generating sentences becomes easier, but the false positive (FP) rate (i.e., human-written text misclassified as machine-generated) increases.
Furthermore, the choice of interval affects the robustness of SimMark against paraphrasing attacks. When an attacker paraphrases sentences, the similarities (or distances) may change and fall outside the interval. A larger interval provides greater robustness, as the watermark is less likely to be disrupted by paraphrasing. Therefore, the selection of the interval involves balancing several factors: it must not be too narrow to impede sentence generation while maintaining a low FP rate and adequate robustness against paraphrasing.
Finding an optimal interval, or “sweet spot,” depends on the distribution of similarities between consecutive sentences, which varies across different embedding models. However, identifying such sweet spots is feasible when the user analyzes the similarity distributions of both LLM-generated and human-written text (See Appendix E for an example of finding the sweet spot).
3.2 Watermarked Text Detection
The detection of SimMark follows a similar methodology to KGW by employing a one-proportion $z$-test for hypothesis testing. However, akin to Hou et al. (2023, 2024), the detection operates at the sentence level rather than the token level. To perform detection, we first divide the input text into sentences [^2] Using sent_tokenize method of NLTK (Bird et al., 2009). and use the same Instructor-Large embedding model with identical instruction to compute the embeddings for each sentence. If PCA was used during the watermark generation process, it must also be applied here to ensure consistency.
Algorithm 2 Detection Pseudo-Code for SimMark
0:  input text, embedding model, predefined interval $[a,b]$, decay factor $K$, threshold $\beta$, optional PCA model
1:  Split the input text into sentences excluding the first sentence (i.e, the prompt): $M=\{X_{2},\dots,X_{N_{\mathrm{total}}}\}$.
2:  for each sentence pair $(X_{i},X_{i+1})$ do
3:     Compute embeddings $e_{i}$ for $X_{i}$ and $e_{i+1}$ for $X_{i+1}$ using the embedding model.
4:     Optional: Reduce the dimensionality of $e_{i}$ and $e_{i+1}$ using the PCA model.
5:     Compute $s_{i+1}\leftarrow sim(e_{i},e_{i+1})$.
6:     Compute $c_{i+1}$ according to Eq. (1).
7:  end for
8:  $N\leftarrow|M|$
9:  Soft count of valid sentences: $N_{\mathrm{valid\_soft}}\leftarrow\sum_{i=2}^{N+1}c_{i}$.
10:  Estimate $p_{0}$ as the area under the human-written text embeddings similarity distribution curve within $[a,b]$.
11:  Compute $z_{\mathrm{soft}}$ (i.e., the soft-$z$-score) using Eq. (2).
12:  if $z_{\mathrm{soft}}>\beta$ then
13:     Reject $H_{0}$, i.e., the input text is likely generated by an LLM watermarked by SimMark.  [Uncaptioned image]
14:  else
15:     Accept $H_{0}$, i.e., the input text is likely written by a human.   [Uncaptioned image]
16:  end if
Next, we compute the similarity (or distance) between consecutive sentences and count the number of valid sentences, denoted as $N_{\mathrm{valid\_soft}}$. A sentence $X_{i+1}$ is deemed valid if $s_{i+1}$ the similarity (or distance) between $X_{i+1}$ and $X_{i}$ lies within the predefined user interval $[a,b]$. However, paraphrasing attacks may alter embeddings significantly, causing the similarity (or distance) to deviate from the desired interval. To mitigate this, we adopt a soft counting approach, where a sentence is considered partially valid if its similarity is near the interval boundaries. Specifically, the soft count of $X_{i+1}$, denoted as $c_{i+1}$, is defined as follows:
$c_{i+1}=c(s_{i+1})=\begin{cases}1&\text{if }s_{i+1}\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\in[a,b]},\\ e^{-K\min\{|a-s_{i+1}|,|b-s_{i+1}|\}}&{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\text{otherwise.}}\end{cases}\end{aligned}$
Here, $s_{i+1}=sim(e_{i},e_{i+1})$ represents the similarity (or distance) between consecutive sentence embeddings, and $K>0$ is a decay factor controlling the smoothness of the soft counting mechanism. A higher $K$ makes the function behave closer to a step function, while a lower $K$ allows for smoother transitions, tolerating minor deviations outside the interval $[a,b]$. The total number of valid sentences is then computed as $N_{\mathrm{valid\_soft}}=\sum c_{i}$.
Refer to Appendix C for experimental results demonstrating how this soft counting mechanism improves robustness against paraphrasing, with only a minimal impact on performance in non-paraphrased scenarios. Allowing for some degree of error through soft counting enables better detection under adversarial conditions while maintaining high accuracy in non-paraphrased cases.
During our initial experiments, we observed that contrary to cosine similarity, Euclidean distance is very sensitive to paraphrasing and even a subtle change in the sentences would result in a huge difference in the distances of embeddings. We hypothesize that Euclidean distance is sensitive to noise in high-dimensional spaces such as the semantic space of the embedder. To mitigate this, we propose using PCA: We first train a PCA model on a dataset of human-written texts, to find the principal components of the sentence embeddings, i.e., the components that contribute the most to the semantical representation of the sentences. Then, we apply PCA to the embeddings to reduce their dimension. More details on the dimensionality reduction are provided in Section 4 and Appendix D.
With this approach reverse-engineering the watermark algorithm becomes even harder as the third party who tries to reverse-engineer the algorithm, not only requires to know the embedder model and the instruction used for it, but it also needs to know the specifics of PCA setting (such as the number of principal components) and needs access to the dataset for training the PCA; otherwise, the similarity distribution of sentence embeddings would be different, making reverse engineering challenging. We found that this approach improved the robustness against paraphrasing when Euclidean distance is used. However, it would decrease the performance when cosine similarity is used.
Returning to the detection, the null hypothesis $H_{0}$ is defined as follows:
$H_{0}$: The sentences are written by humans, i.e., the text sequence is written (or generated) without knowledge of the valid interval in the similarity distribution of sentence embeddings.
We calculate the $z$-statistic for the one-proportion $z$-test using the sample proportion $p=\frac{N_{\mathrm{valid\_soft}}}{N}$, where $N$ is the total number of samples (sentences). Since $N_{\mathrm{valid\_soft}}$ is a soft count of valid sentences, we refer to it as soft-$z$-score, which is given by $z_{\mathrm{soft}}=\frac{p-p_{0}}{\sqrt{\frac{p_{0}(1-p_{0})}{N}}}$. Here, $p_{0}$ is the population proportion and represents the ratio of valid sentences to all sentences in human-written text (i.e., in a text with no watermark), which is estimated as the area under the similarity distribution curve of consecutive human-written sentences within the interval $[a,b]$ (Like the one in Figure 6). Alternatively, the soft-$z$-score can be expressed as:
$z_{\mathrm{soft}}=\tfrac{N_{\mathrm{valid\_soft}}-p_{0}N}{\sqrt{p_{0}(1-p_{0})N}}.$ (2)
In this form, the value of $z_{\mathrm{soft}}$ can be interpreted as a normalized deviation of the number of valid sentences $N_{\mathrm{valid\_soft}}$ from its expectation $p_{0}N$. As highlighted in Algorithm 2, the null hypothesis $H_{0}$ is rejected if $z_{soft}>\beta$, where $\beta$ is a threshold determined empirically by running the detection algorithm on human-written text. The threshold $\beta$ is selected to maintain a desired FP rate (i.e., minimizing the misclassification of human-written text as LLM-generated). Details on the computation of $\beta$ are provided in the Appendix F.
4 Experiments
Dataset Algorithm No Paraphrase Pegasus Pegasus-bigram Parrot Parrot-bigram GPT3.5 GPT3.5-bigram Avg. Paraphrased
RealNews KGW (Kirchenbauer et al.) 99.6 / 98.4 / 98.9 95.9 / 82.1 / 91.0 92.1 / 42.7 / 72.9 88.5 / 31.5 / 55.4 83.0 / 15.0 / 39.9 82.8 / 17.4 / 46.7 75.1 / 5.9 / 26.3 86.2 / 32.4 / 55.4
SIR (Liu et al.) 99.9 / 99.4 / 99.9 94.4 / 79.2 / 85.4 94.1 / 72.6 / 82.6 93.2 / 62.8 / 75.9 95.2 / 66.4 / 80.2 80.2 / 24.7 / 42.7 77.7 / 20.9 / 36.4 89.1 / 54.4 / 67.2
SemStamp (Hou et al.) 99.2 / 93.9 / 97.1 97.8 / 83.7 / 92.0 96.5 / 76.7 / 86.8 93.3 / 56.2 / 75.5 93.1 / 54.4 / 74.0 83.3 / 33.9 / 52.9 82.2 / 31.3 / 48.7 91.0 / 56.0 / 71.6
$k$-SemStamp (Hou et al.) 99.6 / 98.1 / 98.7 99.5 / 92.7 / 96.5 99.0 / 88.4 / 94.3 97.8 / 78.7 / 89.4 97.5 / 78.3 / 87.3 90.8 / 55.5 / 71.8 88.9 / 50.2 / 66.1 95.6 / 74.0 / 84.2
Cosine-SimMark (Ours) 99.6 / 96.8 / 98.8 99.2 / 90.3 / 98.2 99.1 / 90.3 / 97.9 98.7 / 88.1 / 97.2 98.8 / 87.3 / 97.6 95.7 / 59.7 / 86.7 92.0 / 38.8 / 73.7 97.2 / 75.8 / 91.9
Euclidean-SimMark [^7] PCA is applied. (Ours) 99.8 / 98.5 / 99.3 97.2 / 72.3 / 89.1 96.9 / 70.0 / 87.4 95.7 / 60.2 / 82.5 95.7 / 59.1 / 81.5 94.1 / 51.6 / 76.2 88.2 / 29.7 / 53.5 94.6 / 57.2 / 78.4
BookSum KGW 99.6 / 99.0 / 99.2 97.3 / 89.7 / 95.3 96.5 / 56.6 / 85.3 94.6 / 42.0 / 75.8 93.1 / 37.4 / 71.2 87.6 / 17.2 / 52.1 77.1 / 4.4 / 27.1 91.0 / 41.2 / 67.8
SIR 100 / 99.8 / 100 93.1 / 79.3 / 85.9 93.7 / 69.9 / 81.5 96.5 / 72.9 / 85.1 97.2 / 76.5 / 88.0 80.9 / 39.9 / 23.6 75.8 / 19.9 / 35.4 89.5 / 59.7 / 66.6
SemStamp 99.6 / 98.3 / 98.8 99.0 / 94.3 / 97.0 98.6 / 90.6 / 95.5 98.3 / 83.0 / 91.5 98.4 / 85.7 / 92.5 89.6 / 45.6 / 62.4 86.2 / 37.4 / 53.8 95.0 / 72.8 / 82.1
$k$-SemStamp 99.9 / 99.1 / 99.4 99.3 / 94.1 / 97.3 99.1 / 92.5 / 96.9 98.4 / 86.3 / 93.9 98.8 / 88.9 / 94.9 95.6 / 65.7 / 83.0 95.7 / 64.5 / 81.4 97.8 / 81.5 / 91.2
Cosine-SimMark (Ours) 99.8 / 98.8 / 99.5 99.5 / 93.3 / 98.5 99.6 / 94.1 / 98.5 99.3 / 88.5 / 98.0 99.3 / 87.0 / 98.2 97.1 / 62.5 / 86.9 94.5 / 41.6 / 74.2 98.2 / 77.8 / 92.4
Euclidean-SimMark (Ours) 100 / 100 / 100 98.8 / 82.6 / 94.9 98.6 / 80.4 / 93.4 97.9 / 75.3 / 91.1 97.9 / 73.3 / 91.6 99.7 / 94.4 / 98.8 99.5 / 91.9 / 97.6 98.7 / 83.0 / 94.6
Reddit-TIFU KGW 99.3 / 97.5 / 98.1 94.1 / 87.2 / 87.2 91.7 / 67.2 / 67.6 79.5 / 22.8 / 43.3 82.8 / 27.6 / 49.7 84.1 / 27.3 / 50.9 79.8 / 19.3 / 41.3 85.3 / 41.9 / 56.7
SemStamp 99.7 / 97.7 / 98.2 98.4 / 92.8 / 95.4 98.0 / 89.0 / 92.9 90.2 / 56.2 / 70.5 93.9 / 71.8 / 82.3 87.7 / 47.5 / 58.2 87.4 / 43.8 / 55.9 92.6 / 66.9 / 75.9
Cosine-SimMark (Ours) 99.1 / 96.3 / 97.6 98.9 / 94.5 / 96.4 98.7 / 93.6 / 96.1 98.5 / 91.6 / 96.0 98.5 / 91.7 / 95.5 97.8 / 88.4 / 94.7 96.3 / 72.9 / 88.4 98.1 / 88.8 / 94.5
Euclidean-SimMark [^7] PCA is applied. (Ours) 99.8 / 98.7 / 99.2 99.0 / 94.7 / 97.6 99.0 / 91.9 / 96.2 97.8 / 75.9 / 89.5 97.7 / 76.4 / 90.4 98.7 / 83.7 / 95.2 96.8 / 65.8 / 87.3 98.2 / 81.4 / 92.7
[…]
No watermark 11.89 11.43 3.10
SemStamp 12.89 11.51 3.17
$k$-SemStamp 11.82 11.48 3.11
Cosine-SimMark (Ours) 12.69 11.50 3.39
Euclidean-SimMark [^7] PCA is applied. (Ours) 9.67 11.50 3.28
Table 2: Comparison of the quality of text generated using different watermarking algorithms on the BookSum dataset. For PPL metric, lower values indicate better quality, while higher values are better for Ent-3 and Sem-Ent. SimMark watermarking achieves quality comparable to no watermarking, demonstrating minimal impact on text quality.
In this section, we evaluate the performance of our watermarking algorithm, SimMark, across different datasets. Performance is assessed using the area under the receiver operating characteristic curve (ROC-AUC) and true positive rate (TP) at fixed false positive rates (FP) of 1% and 5%, (TP@1%FP and TP@5%FP). Higher values indicate better performance across all metrics.
For dimensionality reduction, we trained a PCA model on 8000 samples from the RealNews subset of the C4 dataset (Raffel et al., 2020), reducing embedding dimensions from 768 to 16. After testing various principal component counts (ranging from 512 to 16), we found 16 to yield the best results. PCA improved robustness against paraphrasing attacks when Euclidean distance was used (except on the BookSum dataset) but consistently degraded performance when cosine similarity was employed across all datasets. During our experiments, we applied PCA wherever it enhanced performance. More details and the results of these experiments are summarized in Table 4 in Appendix D.
Across all experiments, the decay factor was set to $K=250$, as this value provided an optimal trade-off between performance under both non-paraphrased and paraphrased conditions (See Appendix C). The threshold $\beta$ was determined programmatically during the detection (See Appendix F) to achieve the specified FP rates (1% or 5%). The predefined intervals $[0.68,0.76]$ for cosine similarity and $[0.28,0.36]$ for Euclidean distance with PCA and $[0.4,0.55]$ for Euclidean distance without PCA were found to be near-optimal. While the intervals could have been further optimized for each dataset individually, we chose not to do so to demonstrate the general performance of our method. For additional details on the experimental environment refer to Appendix H.
4.1 Models and Datasets
To ensure comparability with prior works by Hou et al. (2023, 2024), we also employed the same fine-tuned version of OPT-1.3B (Zhang et al., 2022), that they used as our LLM [^3] Used AbeHou/opt-1.3b-semstamp model.. For semantic embedding, we utilized the Instructor-Large model [^4] Used hkunlp/instructor-large embedding model. (Su et al., 2023), as mentioned earlier. Following Kirchenbauer et al. (2023), sampling from the LLM was performed with a temperature of 0.7 and a repetition penalty of 1.05, while the minimum and the maximum number of generated tokens were set to 195 and 205, respectively. The maximum number of iterations for rejection sampling is set to 100, following Hou et al. (2023, 2024).
In our experiments, we used the RealNews subset of the C4, the BookSum (Kryscinski et al., 2022), and Reddit-TIFU (Kim et al., 2019) datasets, as in Hou et al. (2023). 1000 samples from each dataset were chosen to analyze the detection results and the text quality. Each sample was segmented into sentences, with the first sentence serving as the prompt to the LLM.
We evaluated text quality after applying SimMark using the following metrics:
  • Perplexity (PPL): Measures how surprising the text is to a larger LLM (OPT-2.7B (Zhang et al., 2022)). Lower values indicate higher quality.
  • Tri-gram Entropy (Ent-3) (Zhang et al., 2018): To assess textual diversity via the entropy of the tri-grams distribution.
  • Semantic Entropy (Sem-Ent) (Han et al., 2022): Used to measure semantic informativeness and diversity of the text.
4.2 Paraphrasing Attack
To evaluate the robustness of SimMark, we tested it against paraphrasing attacks using three models: Pegasus paraphraser [^5] Used tuner007/pegasus_paraphrase paraphrasing model. (Zhang et al., 2020), Parrot paraphraser (Damodaran, 2021), and GPT-3.5-Turbo (OpenAI, 2022). Kirchenbauer et al. (2024) observed that prompting models to paraphrase entire texts often results in summarized outputs, with the summarization ratio worsening for longer inputs. To prevent such summarization, we adopted a sentence-by-sentence paraphrasing scheme, which also ensures our results are comparable to prior works such as Hou et al. (2023, 2024). The quality of paraphrases was assessed using BertScore [^6] Used microsoft/deberta-xlarge-mnli (He et al., 2021) model. (Zhang* et al., 2020), with all settings consistent with prior works of Hou et al. (2023, 2024). The bottom part of Figure 2 demonstrates the paraphrase attack and detection phase in more detail. Refer to Appendix I to find the prompts used with GPT-3.5-Turbo.
We also included the results for the bigram paraphrase attacks introduced by Hou et al. (2023), with identical settings (25 rephrases for each sentence, etc.). This type of attack involves generating multiple paraphrases for each sentence, and choosing the one that increases the likelihood of disrupting statistical signatures embedded in the text, especially for token-level watermarking algorithms (Hou et al., 2023). While this type of attack significantly impacts most other models, SimMark demonstrates greater robustness against it. We must highlight that direct comparisons to $k$-SemStamp are not entirely fair, as $k$-SemStamp relies on domain-specific clustering of semantic spaces, making it domain-dependent. In contrast, SimMark, similar to SemStamp, is domain-independent. Still, SimMark outperforms $k$-SemStamp in most cases across various datasets and metrics, further underscoring its universality and robustness.
4.3 Paraphrase+Drop Attack
Figure 3: Detection results (ROC-AUC) under the Paraphrase+Drop Attack. Pegasus is used as the paraphraser. Here, 0 probability means no sentence is dropped, i.e., sentences are only paraphrased. SimMark is more robust than SemStamp and $k$-SemStamp across varying sentence drop probabilities.
To further evaluate the robustness of our method and its real-world applicability, we introduce a more challenging attack scenario, termed the Paraphrase+Drop Attack. We argue that an adversary not only paraphrases the text but also drops sentences deemed redundant. This reflects common editing practices since LLMs are known for their tendency to generate lengthy outputs with redundancy, likely due to their imperfect reward modelling (Chiang & Lee, 2024).
To simulate this scenario, we first paraphrase the text using the method described in the previous subsection and then randomly drop sentences from the paraphrased text. This combined attack tests the resilience of SimMark under more aggressive transformations that mimic real-world adversarial editing.
4.4 Results & Discussion
We compared the performance of SimMark against state-of-the-art watermarking algorithms through extensive experiments. Our primary baseline was SemStamp, a sentence-level semantic watermarking method. We also included $k$-SemStamp, an improvement over SemStamp tailored to specific text domains. Results for SemStamp, $k$-SemStamp, and other methods, including KGW and SIR, were extracted directly from Hou et al. (2024).
Table 1 presents detection performance before and after paraphrasing attacks, while Table 2 provides text quality evaluation results. Our algorithm (the last two rows in each section) does not affect the text quality while being effective and consistently outperforming or matching other state-of-the-art methods, achieving the highest average performance across all paraphrasers and datasets. Notably, domain-independent SimMark surpasses our primary baseline, SemStamp, and matches or exceeds the domain-dependent $k$-SemStamp. Refer to Appendix B for qualitative examples of SimMark.
A key aspect to consider is that the fine-tuning of SemStamp and $k$-SemStamp’s embedding model on text paraphrased by Pegasus likely contributes to its improved robustness against this specific paraphraser but may introduce bias. Additionally, the results for the Reddit-TIFU dataset were only available for SemStamp and not $k$-SemStamp, likely due to the dataset’s informal diverse text style and $k$-SemStamp’s limitation for text to belong to a specific domain such as news articles or scientific writings (Hou et al., 2024).
Figure 3 presents the detection performance (in terms of ROC-AUC) of SimMark, $k$-SemStamp, and SemStamp against Paraphrase+Delete Attack under varying sentence drop probabilities, evaluated on the RealNews and BookSum datasets. The results demonstrate that SimMark consistently achieves higher performance compared to other methods, even as the sentence drop probability increases. These findings underscore the robustness of SimMark in handling more challenging real-world adversarial scenarios.
Despite Hou et al. (2023, 2024) releasing their code and data, we were unable to reproduce their reported results fully. Consequently, there are minor discrepancies between our reproduction results (shown in Figure 3 for the case with zero probability) and those presented in Table 1, which were extracted directly from their paper. Notably, the results for $k$-SemStamp on the BookSum dataset differed substantially, leading us to exclude them from Figure 3. Additionally, the results reported in Table 2 (our reproduction) also show slight differences from their paper, likely due to hyperparameter details that were not explicitly documented.
Regarding sampling efficiency, for BookSum dataset for instance, cosine-SimMark required an average of 7.1 samples per sentence from the LLM, compared to $k$-SemStamp and SemStamp, which averaged 13.3 and 20.9 samples, respectively (Hou et al., 2024). This demonstrates that our method not only outperforms these baselines but is also significantly (about 2-3 times) more efficient. Further analysis of this, including theoretical estimates is provided in Appendix G.
5 Conclusion
In this paper, we introduced SimMark, a similarity-based, sentence-level posthoc watermarking algorithm. Unlike existing approaches, SimMark operates without requiring access to the internal logits or architecture of the model, ensuring compatibility with a wide range of LLMs, including API-only models. By utilizing a general-purpose sentence embedding model and integrating a soft counting mechanism, SimMark combines robustness against paraphrasing with applicability to diverse domains. Furthermore, its resistance to reverse engineering ensures resilience against spoofing attacks. Experimental results show that SimMark outperforms state-of-the-art watermarking algorithms, such as SemStamp, in both efficiency and robustness to adversarial paraphrasing, representing a significant step forward in practical watermarking techniques for LLM-generated content. Due to space constraints, the discussion of the future directions is deferred to Appendix A.
Acknowledgements
This work was supported by the NSERC Discovery Grant No. RGPIN-2019-05448.
Impact Statement [^7] Generative AI tools, such as ChatGPT (OpenAI, 2022), were used to refine this manuscript. The authors retain full responsibility for all content presented in the paper, ensuring adherence to academic integrity and ethical research practices.
By enabling robust detection of LLM-generated text, particularly under paraphrasing attacks, SimMark addresses ethical concerns surrounding the transparency and accountability of AI-generated content. However, like any watermarking method, there are potential risks, such as falsely implicating human authors or adversaries developing more advanced techniques for spoofing attacks or bypassing detection.
We acknowledge these limitations and advocate for the responsible deployment of such tools in combination with other verification mechanisms to mitigate these risks and ensure ethical, fair deployment. The primary goal of this work is to advance research in watermarking techniques to support the responsible use of LLMs. We believe that the societal impacts and ethical considerations of our work align with those outlined in Weidinger et al. (2021).
Appendix A Future Work
While SimMark demonstrates outstanding performance, there are still some areas that warrant further exploration:
Reduce Rejection Sampling Overhead. The rejection sampling process requires generating multiple candidate sentences until a valid sentence is accepted. Although our method is 2-3 times more efficient than prior approaches such as SemStamp and $k$-SemStamp, there is still a notable decrease in generation speed due to rejection sampling. Techniques like batch sampling or parallel sampling could potentially mitigate this issue, though at the expense of higher computational resource usage. Future research should focus on optimizing these methods to balance efficiency and resource requirements.
Improve Resistance to More Advanced Attacks. Although SimMark is robust against paraphrasing attacks to some extent, more advanced text transformations (e.g., sentence splitting and merging, semantic reordering) may still pose challenges. Additionally, our detection algorithm may struggle when watermarked text is mixed into a larger body of human-written content. Further research is needed to enhance resilience against a broader range of adversarial transformations.
Setting the Interval Dynamically. In our experiments, we used consistent predefined intervals across all datasets and observed consistently strong performance. Notably, we did not observe any degradation in text quality due to this interval constraint during rejection sampling (as shown in Table 2), likely because the constraint applies only to consecutive sentences. Nonetheless, slight variations in the embeddings similarity distribution of LLM-generated text across different models and domains may impact watermarking effectiveness. Adaptive strategies for setting these intervals dynamically (or pseudo-randomly) could not only improve performance but also make reverse-engineering the algorithm even more difficult.
Appendix B Examples of Watermarked Text
Figures 4 and 5 provide examples of text generated with and without the SimMark watermark. These examples illustrate the imperceptibility of the watermark to human readers while enabling robust detection through our proposed algorithm. They also highlight SimMark’s robustness to paraphrasing while maintaining quality comparable to non-watermarked text.
Figure 4: Example of text generated with and without cosine-SimMark using the RealNews dataset. The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid. Numbers in parentheses represent the soft count for partially valid sentences. The top panel shows non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=0.14\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@cmyk@stroke{0}{1}{1}{0}\pgfsys@color@cmyk@fill{0}{1}{1}{0}<}\;5.033$, false negative). The middle panel demonstrates text generated using SimMark with cosine similarity-based watermarking, producing a strong detection signal ($z_{\text{soft}}=9.48\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;5.033$). The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite semantic alterations ($z_{\text{soft}}=6.94\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;5.033$).
Figure 5: Example of text generated with and without Euclidean-SimMark using the BookSum dataset. The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid. Numbers in parentheses represent the soft count for partially valid sentences. The top panel shows the non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=-1.07\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@cmyk@stroke{0}{1}{1}{0}\pgfsys@color@cmyk@fill{0}{1}{1}{0}<}\;4.13$, false negative). The middle panel demonstrates text generated using SimMark with Euclidean distance-based watermarking, producing a strong detection signal ($z_{\text{soft}}=13.07\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;4.13$). The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite semantic alterations ($z_{\text{soft}}=11.99\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;4.13$).
Appendix C Ablation Study on Soft Count Smoothness Factor $K$
Count Method K cosine-SimMark Paraphrased cosine-SimMark Euclidean-SimMark Paraphrased Euclidean-SimMark
Soft Count 50 99.0 / 89.2 / 97.2 98.6 / 78.5 / 96.5 99.4 / 91.2 / 98.1 96.9 / 49.3 / 87.9
150 99.6 / 95.7 / 98.8 99.2 / 88.7 / 98.2 99.8 / 97.6 / 99.3 97.3 / 67.8 / 90.4
250 99.7 / 96.9 / 98.8 99.2 / 90.3 / 98.2 99.8 / 98.5 / 99.2 97.2 / 72.3 / 88.9
350 99.7 / 96.9 / 98.9 99.2 / 90.4 / 98.1 99.8 / 98.5 / 99.4 97.2 / 71.1 / 88.9
Regular Count $\infty$ 99.7 / 97.2 / 99.1 99.1 / 88.7 / 97.6 99.8 / 98.5 / 99.7 97.0 / 70.0 / 88.2
[…]
In this section, we analyze the impact of the smoothness factor $K$ on the performance of SimMark. Recall that $K$ controls the degree of smoothness in the soft counting mechanism as defined in Eq. (1). A larger $K$ makes the soft counting function behave more like a step function, while smaller values provide smoother transitions between valid and invalid sentences.
Table 3 presents the results of this ablation study, conducted on the RealNews dataset with Pegasus as the paraphraser. Metrics include ROC-AUC, TP@FP=1%, and TP@FP=5%. Higher values indicate better performance across all metrics. The results demonstrate the following trends:
  • A smoothness factor of $K=250$ provides a good trade-off, achieving strong performance both before and after paraphrasing attacks for both cosine-SimMark and Euclidean-SimMark.
  • For $K=\infty$, corresponding to regular counting with a step function, the performance is slightly higher in the absence of paraphrasing but significantly degrades under paraphrasing attacks, highlighting the benefits of soft counting in adversarial scenarios.
These findings confirm that soft counting loses a small amount of performance when no paraphrasing is applied, but it gains substantial robustness under paraphrasing. For example, TP@FP=1% improves by 1.6–2.3% for Pegasus-paraphrased text when $K=250$, and the improvement is likely to be even more significant for stronger paraphrasers.
Appendix D Ablation Study on Impact of PCA
Table 4 presents the results of an ablation study investigating the impact of applying PCA to reduce the dimensionality of sentence embeddings across the RealNews, BookSum, and Reddit-TIFU datasets. Metrics include ROC-AUC, and TP at fixed FP rates (FP=1% and FP=5%). Higher values indicate better performance across all metrics, with PCA applied to embeddings to explore its effect on detection accuracy and robustness.
The results reveal that the effect of PCA depends on the choice of similarity measure. For Euclidean distance-based SimMark, applying PCA generally improves robustness against paraphrasing attacks across most datasets, except for the BookSum dataset. This improvement likely arises because reducing dimensionality helps mitigate noise in the embeddings, especially after paraphrase attack.
On the other hand, for cosine similarity-based SimMark, applying PCA reduces performance across all datasets. This reduction may be due to PCA altering the embeddings in a way that disrupts the angular relationships critical for cosine similarity calculations. These findings highlight the importance of adapting PCA usage based on the similarity measure employed to achieve optimal watermarking performance.
Dataset Configuration No paraphrase Pegasus
RealNews Cosine-SimMark (No PCA) 99.7 / 96.9 / 98.8 99.2 / 90.3 / 98.2
Cosine-SimMark (PCA) 99.6 / 96.9 / 99.1 92.1 / 33.8 / 71.2
Euclidean-SimMark (No PCA) 99.4 / 92.6 / 98.4 90.5 / 19.7 / 58.0
Euclidean-SimMark (PCA) 99.8 / 98.5 / 99.2 97.2 / 72.3 / 88.9
BookSum Cosine-SimMark (No PCA) 99.8 / 98.8 / 99.5 99.5 / 93.3 / 98.5
Cosine-SimMark (PCA) 100 / 99.9 / 99.9 98.7 / 87.3 / 95.1
Euclidean-SimMark (No PCA) 100 / 100 / 100 98.8 / 82.6 / 94.9
Euclidean-SimMark (PCA) 99.9 / 99.3 / 99.5 97.4 / 69.8 / 88.6
Reddit-TIFU Cosine-SimMark (No PCA) 99.1 / 96.3 / 97.6 98.9 / 94.5 / 96.4
Cosine-SimMark (PCA) 99.7 / 98.8 / 99.3 96.6 / 78.9 / 89.3
Euclidean-SimMark (No PCA) 99.6 / 98.1 / 99.1 96.7 / 72.6 / 90.0
Euclidean-SimMark (PCA) 99.8 / 98.7 / 99.2 99.0 / 94.7 / 97.6
[…]
Appendix E Finding an Optimal Interval
Figure 6 shows the distribution of distances between embeddings of consecutive sentences for both human and LLM-generated text, calculated on a sample of size 1000 from the BookSum dataset (no PCA applied to the embeddings in this case). A small but noticeable distribution shift between the two can be observed. Based on this, the interval $[0.4,0.55]$ appears to be a reasonable choice for SimMark watermarking in this case.
It is important to note that changes to the embedding representations, such as applying PCA or using a different embedding model, will lead to altered distance distributions. Consequently, the interval must be adjusted accordingly to maintain optimal performance. For instance, if PCA is applied, the interval $[0.28,0.36]$ is suitable. Similarly, if we plot the figure for when cosine similarity is used instead of Euclidean distance, intervals $[0.81,0.94]$ and $[0.68,0.76]$ are good candidates for cases with and without PCA, respectively. This variability in the distance distributions also strengthens the algorithm’s resistance to reverse engineering.
Selecting the optimal interval $[a,b]$ is a critical step in achieving a robust and reliable watermarking with SimMark. In general, selecting an optimal interval involves balancing low FP rates, high TP rates, and robustness against paraphrasing attacks. It is often beneficial to choose intervals toward the tails of the distribution rather than around the mean. Finally, further exploration of dynamic interval selection mechanisms could enhance SimMark’s robustness.
Figure 6: Distribution of Euclidean distances between embeddings of consecutive sentences for human-written and LLM-generated text on the BookSum dataset, generated using OPT-1.3B. The figure demonstrates that the interval $[0.4,0.55]$ is a reasonable choice for Euclidean-SimMark in this case, though it is not necessarily the only viable option.
Appendix F Computing Threshold $\beta$ for soft-$z$-test
Recall that a text is classified as LLM-generated when $z_{\mathrm{soft}}>\beta,$ and as human-written otherwise. $z_{\mathrm{soft}}$ is the statistic used in the statistical test described in Eq. (2). To determine the threshold $\beta$ that limits the FP rate to 1% or 5%, we first need to estimate $p_{0}$, the probability that the consecutive embeddings’ similarity or distance falls within the predefined interval $[a,b]$. This value of $p_{0}$ is a key component in calculating the $z_{\mathrm{soft}}$, as it represents the proportion of valid sentences in human-written text under the given interval. $p_{0}$ serves as an indicator of how frequently valid sentences are expected to occur in human-written text.
To compute $p_{0}$, we analyze the distribution of similarities (or distances) using a histogram approach, such as the one depicted in Figure 6. Specifically, we employ a binning technique to approximate the area under the curve of distribution in the interval $[a,b]$. The process involves dividing the entire range of distances or similarities into a fixed number of bins—1000 bins in our implementation. Each bin represents a small segment of the range, and the histogram is used to calculate the proportion of samples that fall within the interval $[a,b]$.
Mathematically, $p_{0}$ is estimated as:
$\small p_{0}=\frac{\text{Number of samples in bins corresponding to }[a,b]}{\text{Size of the dataset}},$
Once $p_{0}$ is estimated, the detection threshold $\beta$ is determined by iterating over a range of possible values, typically from -10 to 10, to find the one that results in the desired false positive rate. Specifically, the threshold is chosen such that the proportion of human-written texts misclassified as LLM-generated matches the target FP rate (e.g., 1% or 5%).
It is worth noting that the distribution of similarities or distances may vary depending on different factors such as embedding model, and similarity measure (e.g., cosine or Euclidean). As a result, $p_{0}$ and therefore $\beta$ are determined programmatically during the detection to ensure reliable performance of the watermarking algorithm.
Appendix G Analysis of Sampling Efficiency
The average number of samples required to generate a valid sentence is influenced by the chosen interval $[a,b]$. To provide further insights into this relationship, we estimated the area under the curve (AUC) of the embedding similarity distribution for an unwatermarked LLM. Specifically, for the interval $[0.68,0.76]$ (for cosine-SimMark), the estimated AUC is approximately $0.194$.
Using the mean of the geometric distribution, which is given by $\frac{1}{p}$, where $p$ is the probability of success (in this case, the probability of falling within the interval), this translates to an expected average of $\frac{1}{0.194}\approx 5.1$ samples per valid sentence. This estimate is on par with the experimental results observed in our study. The AUC estimates were computed using the binning technique, as described in Appendix F.
This analysis underscores the importance of carefully selecting the interval $[a,b]$, as narrower intervals may increase the number of required samples, leading to reduced sampling efficiency but better performance, while broader intervals may compromise the effectiveness of the watermark. By understanding the interplay between the interval choice and sampling efficiency, we can better optimize SimMark’s performance for different embedding models.
Appendix H Experimental Environment
The majority of the experiments, including text generation and detection tasks, were conducted on a workstation equipped with an Intel Core i9 processor, 64GB of RAM, and an Nvidia RTX 3090 GPU with 24GB of VRAM. Some of the experiments involving bigram paraphrasing were performed on compute nodes with Nvidia V100 GPU with 32GB of VRAM.
Appendix I Prompts Used with GPT-3.5-Turbo
Table 5 presents the prompts we used to obtain paraphrases using GPT-3.5-Turbo (accessed via OpenAI API [^8] https://platform.openai.com/docs/api-reference) for both regular paraphrasing and more aggressive bigram paraphrasing attacks. By using the same prompts as Hou et al. (2023), we ensured that our results were directly comparable to those extracted from their paper, maintaining consistency in evaluation methodology.
Prompt for Regular Attack
[…]
Prompt for Bigram Attack
[…]
Table 5: Prompts used to generate paraphrases with GPT-3.5-Turbo for regular and bigram attacks. These are the same prompts used by Hou et al. (2023) for consistent and comparable evaluation. Here, sent represents the target sentence to rephrase, context includes all preceding sentences, and num_beams specifies the number of paraphrases generated for the bigram attack. A higher num_beams value indicates a more aggressive attack.

### v2

@@
+SimMark: A Robust Sentence-Level Similarity-Based Watermarking Algorithm for Large Language Models
+Abstract
+The widespread adoption of large language models (LLMs) necessitates reliable methods to detect LLM-generated text.
+We introduce SimMark, a robust sentence-level watermarking algorithm that makes LLMs’ outputs traceable without requiring access to model internals, making it compatible with both open and API-based LLMs.
+By leveraging the similarity of semantic sentence embeddings combined with rejection sampling to embed detectable statistical patterns imperceptible to humans, and employing a soft counting mechanism, SimMark achieves robustness against paraphrasing attacks.
+Experimental results demonstrate that SimMark sets a new benchmark for robust watermarking of LLM-generated content, surpassing prior sentence-level watermarking techniques in robustness, sampling efficiency, and applicability across diverse domains, all while maintaining the text quality and fluency.
+[^1] The source code of our algorithm is available here.
+EMNLP’25
+SimMark: A Robust Sentence-Level Similarity-Based
 Watermarking Algorithm for Large Language Models
-Abstract
-The rapid proliferation of large language models (LLMs) has created an urgent need for reliable methods to detect whether a text is generated by such models.
-In this paper, we propose SimMark, a posthoc watermarking algorithm that makes LLMs’ outputs traceable without requiring access to the model’s internal logits, enabling compatibility with a wide range of LLMs, including API-only models.
-By leveraging the similarity of semantic sentence embeddings and rejection sampling to impose detectable statistical patterns imperceptible to humans, and employing a soft counting mechanism, SimMark achieves robustness against paraphrasing attacks.
-Experimental results demonstrate that SimMark sets a new benchmark for robust watermarking of LLM-generated content, surpassing prior sentence-level watermarking techniques in robustness, sampling efficiency, and applicability across diverse domains, all while preserving the text quality.***The source code of our algorithm can be found at https://github.com/DabiriAghdam/SimMark.
-Large Language Models,Watermarking LLMs,Posthoc Watermarking,Sentence-Level Watermarking,Similarity-Based Watermarking, Machine Learning, ICML
+Amirhossein Dabiriaghdam  and Lele Wang Department of ECE, University of British Columbia, Vancouver, BC, Canada {amirhossein, lelewang}@ece.ubc.ca
 1 Introduction
+The advent of deep generative models has made it increasingly important to determine whether a given text, image, or video was produced by artificial intelligence (AI), and recently, researchers across various domains have begun tackling this challenge Aaronson and Kirchner (2022); Fernandez et al.
+(2023); Teymoorianfard et al.
+(2025).
+In particular, LLMs such as GPT-4o Hurst et al.
+(2024), can now generate human-like text at scale and low cost, enabling powerful applications across numerous industries and areas such as health care and law Singhal et al.
+(2023); Wu et al.
+(2025); Torabi et al.
+(2025); Yao et al.
+(2024); Taranukhin et al.
+(2024).
+This capability, however, introduces serious risks, including academic plagiarism, disinformation campaigns, and public opinion manipulation.
+For instance, the use of AI-generated content in news articles has raised concerns about transparency, accountability, and the spread of false information Futurism (2023).
+Moreover, reliably detecting LLM-generated content is crucial for enforcing copyright protections and ensuring accountability Weidinger et al.
+(2021).
 Figure 1: A high-level overview of SimMark detection algorithm.
-The input text is divided into individual sentences $X_{1}$ to $X_{N}$, which are embedded using a semantic embedding model, with optional PCA for dimensionality reduction.
+The input text is divided into individual sentences $X_{1}$ to $X_{N}$, which are embedded using a semantic embedding model.
 The similarity between consecutive sentence embeddings is computed.
 Sentences with similarities within a predefined interval $[a,b]$ are considered valid, while those outside are invalid.
-A soft-$z$-test is performed using the soft count of invalid/partially valid sentences to determine whether the text is watermarked.
-The advent of deep generative AIs has made it increasingly important to detect whether a given text is generated by a large language model (LLM).
-Advanced LLMs, such as GPT-3.5 (OpenAI, 2022), GPT-4 (OpenAI et al., 2023), and Claude 3 (Anthropic, 2024), can produce human-like text at scale and low cost, enabling transformative applications across industries.
-However, these capabilities also present significant risks, including but not limited to misuse for academic plagiarism, disinformation campaigns and propaganda aimed at manipulating public opinion.
-For example, the use of AI-generated content in news articles has raised concerns about transparency, accountability, and the potential spread of misinformation (Futurism, 2023).
-Furthermore, the ability to reliably identify LLM-generated content is essential for enforcing copyright protections and ensuring accountability for such content (Weidinger et al., 2021).
+A statistical test is performed using the count of valid sentences to determine whether the text is watermarked.
 Detecting LLM-generated text poses a unique challenge.
@@
 (2023), reliably differentiating between human-written and machine-generated text remains an open problem.
-One promising approach to addressing this, is the use of imperceptible statistical signatures, or watermarks, embedded within a text.
-Watermarking imperceptibly alters text such that it remains natural to human readers but enables subsequent detection of its origin (Atallah et al., 2001).
-Effective watermarking methods must balance the preservation of LLM output quality with robustness against adversarial paraphrasing, where attackers modify the text to evade detection (Krishna et al., 2024).
-Additionally, watermarks must be resistant to spoofing attacks, wherein adversaries craft non-machine-generated text (often malicious) to falsely trigger watermark detectors (Sadasivan et al., 2023).
-In this paper, we introduce SimMark, a robust posthoc sentence-level watermarking algorithm for LLMs based on sentence embedding similarity.
-The term posthoc here refers to the fact that the watermarking process operates on text after it has been generated by the LLM, rather than modifying the text generation process itself.
-This approach makes SimMark compatible with a wide range of models, including open-weight LLMs and closed-source proprietary models accessible only via APIs, as it does not require fine-tuning or access to the model’s internal logits.
-Access to logits is often restricted by API providers due to their potential use in distilling LLMs and leaking proprietary information (Finlayson et al., 2024).
-SimMark leverages embedding representations from text embedding models to capture semantic relationships between sentences and imposes detectable statistical patterns on sentence similarity through rejection sampling.
-Specifically, rejection sampling involves querying the LLM multiple times until the similarity between embeddings of consecutive sentences falls within a certain predefined interval.
-At detection time, these patterns are analyzed using one-proportion soft-$z$-test, a statistical test, to differentiate between human-written and LLM-generated text, as illustrated in Figure 1.
+One promising approach is the use of imperceptible statistical signatures, or watermarks, embedded within a text.
+Watermarking imperceptibly alters text such that it remains natural to human readers but enables subsequent detection of its origin Atallah et al.
+(2001).
+Effective watermarking must balance the preservation of the text quality with robustness against adversarial paraphrasing, where attackers modify the text to evade detection Krishna et al.
+(2024).
+Additionally, watermarks must be resistant to spoofing attacks, wherein adversaries craft non-machine-generated text (often malicious) to falsely trigger detectors Sadasivan et al.
+(2023).
+In this paper, we introduce SimMark, a robust sentence-level watermarking algorithm for LLMs based on sentence embedding similarity.
+SimMark treats LLMs as black boxes that can be prompted to generate sentences given a context.
+This approach makes SimMark compatible with a wide range of models, including open-weight LLMs and closed-source proprietary models accessible only via APIs, as it does not require fine-tuning or access to the models’ internal logits.
+Access to logits is often restricted by API providers due to their potential use in distilling LLMs and leaking proprietary information Finlayson et al.
+(2024).
+SimMark leverages embeddings from semantic text embedding models to capture semantic relationships between sentences and embeds detectable statistical patterns on sentence similarity through rejection sampling.
+Specifically, rejection sampling involves querying the LLM multiple times until the similarity between the embeddings of consecutive sentences falls within a predefined interval.
+During detection, these patterns are analyzed using a statistical test to differentiate between human-written and LLM-generated text, as illustrated in Figure 1.
 In summary, our contributions are as follows:
-  • We introduce a novel watermarking algorithm that achieves state-of-the-art detection performance while maintaining low false positive rates for human-written text.
+  • We introduce a novel sentence-level watermarking method that achieves state-of-the-art detection performance while maintaining low false positive rates for human-written text.
   • Our approach demonstrates robustness against paraphrasing attacks through semantic-level watermarking and a soft counting mechanism for statistical testing.
-  • Compared to existing watermarking methods, SimMark provides a more practical solution that operates without access to LLM logits or fine-tuning, offering efficient and high-quality watermark injection and detection.
-The remainder of this paper is organized as follows.
-Section 2 reviews the background and related work on watermarking techniques for LLMs.
+  • Compared to existing methods, SimMark provides a more practical solution that operates without access to LLM logits, offering high-quality watermark injection and detection.
+The remainder of this paper is organized as follows: Section 2 reviews the background and related work on LLM watermarking techniques.
 Section 3 outlines our methodology.
 Section 4 describes our experimental setup and presents comparative results, while Section 5 concludes the paper.
-In the end, we also discuss the potential ethical considerations and societal impacts.
 2 Background
@@
 To generate $t_{i+1}$, the LLM samples it from the conditional probability distribution $P(t_{i+1}|T_{i})$ over the vocabulary $V$.
-After generating $t_{i+1}$, the updated sequence $T_{i+1}=\{t_{1},t_{2},\ldots,t_{i},t_{i+1}\}$ is fed back into the model $f$, and the process is repeated iteratively to generate the subsequent token $t_{i+2}$.
-This continues until an end-of-sequence token is generated or the generation reaches a predefined token limit.
-This sequential process of generating one token at a time based on the previously generated tokens is known as autoregressive decoding.
+After generating $t_{i+1}$, the updated sequence $T_{i+1}=T_{i}\cup\{t_{i+1}\}$ is fed back into the model, and the process is repeated iteratively to generate the subsequent tokens.
+This process of generating one token at a time, given the previously generated tokens, is known as autoregressive decoding.
 2.2 Token-Level Watermarking
-Token-level watermarking methods embed a statistical signal in the text by manipulating the token generation process (Kirchenbauer et al., 2023; Zhao et al., 2023; Fu et al., 2024).
-These methods typically alter the probability distribution over $V$ during decoding, subtly biasing the selection of certain tokens to form detectable patterns.
-For instance, the KGW algorithm introduced by Kirchenbauer, Geiping, Wen, Katz, Miers, and Goldstein (2023), groups $V$ into green and red subsets pseudo-randomly seeded on the previous token before generating each new token.
-A predefined positive constant $\delta$ is added to the logits of each token in the green list, increasing their probability of being selected during the sampling step.
-At detection, a one-proportion $z$-test is applied to the number of tokens from the green list in the text to determine whether the text contains a watermark.
-This test compares the observed proportion of green tokens to the expected proportion under the null hypothesis of no watermark, providing a statistical measure to detect even subtle biases introduced by the watermark (Kirchenbauer et al., 2023).
-Zhao et al.
-(2023)’s UNIGRAM-WATERMARK builds upon KGW by fixing the red and green lists instead of pseudo-randomly selecting them, proving that compared to KGW their method is more robust to paraphrasing and editing (Zhao et al., 2023).
-However, as outlined by Hou et al.
-(2024), this algorithm can be reverse-engineered, rendering it impractical for high-stakes, real-world applications.
-The Semantic Invariant Robust (SIR) watermark in Liu et al.
-(2023) is also similar to KGW but is designed to be less sensitive to attacks involving synonym replacement or advanced paraphrasing.
-SIR achieves this by altering the LLM logits based on the semantics of previously generated tokens, using an embedding model to compute semantic representations and training a model that adjusts LLM’s logits based on the semantic embeddings of prior tokens (Liu et al., 2023).
-Detection of such watermarks involves analyzing token sequences for statistical signatures that deviate from typical human-generated text.
-However, token-level watermarks can still be vulnerable to paraphrasing attacks, as rephrasing may disrupt the green and red token lists without altering the overall meaning (Krishna et al., 2024).
-Moreover, since these methods modify the logits, they directly impact the conditional probability distribution over $V$, potentially affecting the quality of the generated text (Fu et al., 2024).
+Token-level watermarking methods embed a statistical signal in the text by manipulating the token sampling process Aaronson and Kirchner (2022); Kirchenbauer et al.
+(2023); Fu et al.
+(2024).
+These methods typically alter the probability distribution over $V$, subtly biasing the selection of certain tokens to form detectable patterns.
+KGW introduced by Kirchenbauer et al.
+(2023), groups $V$ into green and red subsets pseudo-randomly seeded on the previous token before generating each new token.
+A predefined constant $\delta>0$ is added to the logits of each token in the green list, increasing their likelihood of being selected during the sampling step.
+At detection, a $z$-test is applied to the number of tokens from the green list in the text to determine whether the text contains a watermark.
+This test compares the observed proportion of green tokens to the expected proportion under the null hypothesis of no watermark, providing a statistical measure to detect even subtle biases introduced by the watermark.
+Detection of such watermarks involves analyzing tokens for statistical signatures that deviate from typical human text.
+However, token-level watermarks can still be vulnerable to paraphrasing, as rephrasing may disrupt the green and red token lists without altering the overall semantic Krishna et al.
+(2024).
+Moreover, since these methods modify the logits, they directly impact the conditional probability distribution over $V$, potentially degrading the quality of the generated text Fu et al.
+(2024).
+Due to space constraints, additional related work—including token-level methods such as UNIGRAM-WATERMARK (UW) Zhao et al.
+(2023) and the Semantic Invariant Robust (SIR) watermark Liu et al.
+(2023), as well as post-hoc watermarking techniques—is deferred to Appendix A.
 2.3 Sentence-Level Watermarking
-One approach to mitigate the previously mentioned problems is to inject the watermark signal at the sentence-level semantics, making it less vulnerable to paraphrasing attacks.
-Consider a similar notation for sentence generation using model $f$ (i.e., the LLM) with autoregressive decoding that takes a sequence of sentences $M_{i}=\{X_{1},X_{2},\ldots,X_{i}\}$ and generates the next sentence $X_{i+1}$.
+One approach to mitigate the previously mentioned problems is to inject the watermark signal at the sentence level, making it less vulnerable to adversarial modifications Topkara et al.
+(2006).
+Consider a similar notation for sentence generation using an autoregressive LLM that takes a sequence of sentences $M_{i}=\{X_{1},X_{2},\ldots,X_{i}\}$ and generates the next sentence $X_{i+1}$.
 The updated sequence of sentences $M_{i+1}=M_{i}\cup\{X_{i+1}\}$ is then used to generate subsequent sentences iteratively.
-SemStamp (Hou et al., 2023) employs Locality-Sensitive Hashing (LSH) (Indyk & Motwani, 1998) to pseudo-randomly partition the semantic space of the embedding model into a set of valid and blocked regions, analogous to the green and red subsets in KGW.
-During rejection sampling, if the representation of a newly generated sentence lies within the valid regions (determined based on the LSH signature of the previous sentence), the sentence is accepted.
+SemStamp by Hou et al.
+(2024a) employs Locality-Sensitive Hashing (LSH) Indyk and Motwani (1998) to pseudo-randomly partition the semantic space of an embedding model into a set of valid and blocked regions, analogous to the green and red subsets in KGW.
+During rejection sampling, if the embedding of a newly generated sentence lies within the valid regions (determined based on the LSH signature of the previous sentence), the sentence is accepted.
 Otherwise, a new sentence is generated until a valid sentence is produced or the retry limit is reached.
-Similar to KGW, a one-proportion $z$-test is applied to the number of valid sentences (analogous to green tokens) to determine whether the text contains a watermark.
-To improve robustness against paraphrasing, SemStamp used a contrastive learning approach (Hadsell et al., 2006), fine-tuning a semantic embedding model such that the embeddings of paraphrased sentences remain as close as possible to the original sentences.
+Similar to KGW, a $z$-test is applied to the number of valid sentences to determine whether the text contains a watermark.
+To improve robustness against paraphrasing, SemStamp used a contrastive learning approach Hadsell et al.
+(2006), fine-tuning an embedding model such that the embeddings of paraphrased sentences remain as close as possible to the original sentences.
 This was achieved by minimizing the distance between paraphrased and original embeddings while ensuring unrelated sentences remained distinct.
-Additionally, they introduce a margin constraint in the rejection sampling process to reject sentences whose representations lie near the region boundaries.
-Their results demonstrate that this sentence-level approach improves robustness against paraphrasing.
-$k$-SemStamp (Hou et al., 2024) builds upon SemStamp and aims to enhance robustness by partitioning the semantic space using $k$-means clustering (Lloyd, 1982) instead of random partitioning.
-They claim that in this way, sentences with similar semantics are more likely to fall within the same partition, unlike random partitioning, which may place semantically similar sentences into different partitions, reducing robustness (Hou et al., 2024).
-However, $k$-SemStamp assumes that the LLM generates text within a specific domain to apply $k$-means clustering effectively (Hou et al., 2024), limiting its applicability in real-world, open-domain scenarios.
-The generation and detection procedures of $k$-SemStamp remain largely similar to the original SemStamp.
-In contrast to token-level algorithms, these sentence-level methods do not alter the logits of the LLM, therefore it is expected for their resulting text to be of higher quality (Hou et al., 2023).
-Our work, like SemStamp and $k$-SemStamp, is a sentence-level watermarking algorithm; however, it embeds its watermark signature in the semantic similarity of consecutive sentences.
-It achieves greater generalizability across domains by leveraging an off-the-shelf, general-purpose sentence embedding model without fine-tuning.
-At the same time, it outperforms state-of-the-art token-level and sentence-level watermarking methods in robustness against paraphrasing while preserving text quality.
-3 SimMark: A Similarity-Based Watermark
+They also introduce a margin constraint in the rejection sampling process to reject sentences whose embeddings lie near the region boundaries.
+$k$-SemStamp Hou et al.
+(2024b) builds upon SemStamp and aims to enhance robustness by partitioning the semantic space using $k$-means clustering Lloyd (1982) instead of random partitioning.
+They claim that in this way, sentences with similar semantics are more likely to fall within the same partition, unlike random partitioning, which may place semantically similar sentences into different partitions, reducing robustness; however, $k$-SemStamp assumes that the LLM generates text within a specific domain to apply $k$-means clustering effectively Hou et al.
+(2024b), limiting its applicability in real-world, open-domain scenarios.
+The generation and detection procedures of $k$-SemStamp remain similar to the original SemStamp.
+In contrast to token-level algorithms, these sentence-level methods do not alter the internals of the LLM, therefore, it is expected that their output to be of higher quality Hou et al.
+(2024a, b).
+Our work, similar to SemStamp and $k$-SemStamp, is a sentence-level algorithm; however, it injects its watermark signature into the semantic similarity of consecutive sentences.
+It achieves great generalizability across domains by leveraging any off-the-shelf, general-purpose embedding model without fine-tuning.
+At the same time, it outperforms these state-of-the-art (SOTA) sentence-level watermarking methods in robustness against paraphrasing while preserving text quality.
+3 SimMark: A Similarity-Based Watermarking Algorithm
 In this section, we present our proposed framework for watermarking LLMs, detailing both the process of generating watermarked text and its subsequent detection.
-Figure 2: Overview of SimMark algorithm.
+Figure 2: Overview of SimMark.
 Top: Generation.
-For each newly generated sentence ($X_{i+1}$), its embedding ($e_{i+1}$) is computed using the Instructor-Large model, optionally applying PCA for dimensionality reduction.
+For each newly generated sentence ($X_{i+1}$), its embedding ($e_{i+1}$) is computed using a semantic text embedding model, optionally applying PCA for dimensionality reduction.
 The cosine similarity (or Euclidean distance) between $e_{i+1}$ and the embedding of the previous sentence ($e_{i}$), denoted as $s_{i+1}$, is calculated.
@@
 Paraphrased versions of watermarked sentences are generated ($Y_{i}$), and their embeddings ($e^{\prime}_{i}$) are computed.
-The similarity (or distance) between consecutive sentences in the paraphrased text is evaluated.
+The similarity between consecutive sentences in the paraphrased text is evaluated.
 If paraphrasing causes the similarity ($s^{\prime}_{i+1}$) to fall outside $[a,b]$, it is mismarked as invalid.
-A soft counting mechanism (via function $c(s_{i+1})$ instead of a regular counting with a step function in the interval $[a,b]$) quantifies partial validity based on proximity to the interval bounds, enabling detection of watermarked text via the soft-$z$-test even under paraphrase attacks.
-It should be emphasized that soft counting is always applied during detection, regardless of whether paraphrasing is present or not, as we cannot assume prior knowledge of paraphrasing.
+A soft counting mechanism (via function $c(s_{i+1})$ instead of a regular counting with a step function in $[a,b]$) quantifies partial validity based on proximity to the interval bounds, enabling detection of watermarked text via a soft-$z$-test even under paraphrase attacks.
+It should be noted that soft counting is always applied, as we cannot assume prior knowledge of paraphrasing.
 3.1 Watermarked Text Generation
-Similar to SemStamp and $k$-SemStamp (Hou et al., 2023, 2024), SimMark utilizes the embedding representations of the sentences.
+Similar to SemStamp and $k$-SemStamp, SimMark utilizes the embedding representations of the sentences.
 To compute the embeddings, in contrast with Hou et al.
-(2023, 2024) that fine-tuned their embedder model (which could make it biased toward a specific paraphrasing model or domain), we employ Instructor-Large (Su et al., 2023), a general-purpose instruction-tuned embedding model.
-The use of a general-purpose embedding model enables our approach to be more easily adaptable to different domains, and since the generated embeddings are conditioned on a specific instruction, even if a third party knows what embedder model we used, it is still not trivial to reverse-engineer our algorithm without knowing the exact instruction used during generation (the user can also change this instruction frequently), resulting in more robustness against spoofing attacks.
-Algorithm 1 Generation Pseudo-Code for SimMark
-0:  LLM, embedding model, predefined interval $[a,b]$, maximum iterations $N_{\mathrm{max}}$, optional PCA model
-1:  for each generated sentence $i$ ($X_{i}$do
+(2024a, b) that fine-tuned their embedder model (which could make it biased toward a specific paraphrasing model or domain), we employ Instructor-Large Su et al.
+(2023), a general-purpose embedding model, without any fine-tuning.
+The flexibility of our method in using any pretrained embedding model enables our approach to be more easily adaptable to different domains.
+Algorithm 1 SimMark Generation Pseudo-Code
+0: LLM, embedding model, interval $[a,b]$, maximum iterations $N_{\mathrm{max}}$
+1:for each generated sentence $X_{i}$ do
 2:  Compute embedding $e_{i}$ for $X_{i}$ using the embedding model.
@@
 4:  do
-5:         Generate sentence $i+1$ ($X_{i+1}$) using the LLM.
+5:    Generate sentence $X_{i+1}$ using the LLM.
 6:    Compute embedding $e_{i+1}$ for $X_{i+1}$ using the embedding model.
 7:    Optional: Reduce the dimension of $e_{i}$ and $e_{i+1}$ using the PCA model.
-8:         Compute similarity $s_{i+1}\leftarrow sim(e_{i},e_{i+1})$:
-$\small s_{i+1}\leftarrow\begin{cases}\tfrac{e_{i}\cdot e_{i+1}}{\|e_{i}\|_{2}\cdot\|e_{i+1}\|_{2}}&\text{ if cosine similarity},\\ \|e_{i}-e_{i+1}\|_{2}&\text{ if Euclidean distance,}\end{cases}$
+8:    Compute similarity $s_{i+1}$:
+$s_{i+1}\leftarrow\begin{cases}\tfrac{e_{i}\cdot e_{i+1}}{\|e_{i}\|_{2}\cdot\|e_{i+1}\|_{2}}&\text{ if cosine similarity},\\ \|e_{i}-e_{i+1}\|_{2}&\text{ if Euclidean distance,}\end{cases}$
 9:    $n\leftarrow n+1$
-10:     while $s_{i+1}{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\notin[a,b]}$ and $n<N_{\mathrm{max}}$
-11:     Accept $X_{i+1}$ as valid and continue generating the next sentence.
-{Here we either reached the $N_{\mathrm{max}}$ or $s_{i+1}{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\in[a,b]}$}
+10:  while $s_{i+1}{\color[rgb]{1,0,0}\notin[a,b]}$ and $n<N_{\mathrm{max}}$
+11:  Accept $X_{i+1}$ as valid and continue generating the next sentence. {Here we either reached the $N_{\mathrm{max}}$ or $s_{i+1}{\color[rgb]{0,0.5,0}\in[a,b]}$}
 12:end for
-Using Instructor-Large, we compute embeddings for a sentence by passing both the sentence and an instruction (e.g., “Represent the sentence for cosine similarity:” or “Represent the sentence for Euclidean distance:”) to the Instructor model.
-Then, we calculate the cosine similarity (or Euclidean distance) between the embedding of sentence $i+1$ (denoted as $e_{i+1}$ in Algorithm 1) and the embedding of sentence $i$ (denoted as $e_{i}$).
-If the computed value lies within the user-defined predefined interval, sentence $i+1$ is considered valid (analogous to the green subset in KGW).
-Otherwise, we prompt the LLM to generate a new sentence and repeat this procedure until a valid sentence is found or the maximum number of iterations is reached (in this case we accept the last generated sentence, and continue to generating sentence $i+2$), as shown in Algorithm 1.
-Optionally, we may apply Principal Component Analysis (PCA) method to the embeddings to reduce their dimensionality before calculating the similarity or distance (this is the instruction in this case: “Represent the sentence for PCA:”).
+First, we compute the embedding for each sentence [^2] In our experiments, we passed both the sentence and “Represent the sentence for cosine similarity:” or “Represent the sentence for Euclidean distance:” as the instruction to the Instructor-Large model..
+Then, we calculate the cosine similarity (or Euclidean distance) between the embedding of sentence $i+1$ and the embedding of sentence $i$.
+If the computed value lies within a predefined interval, sentence $i+1$ is considered valid (analogous to the green subset in KGW).
+Otherwise, we prompt the LLM to generate a new sentence and repeat this procedure until a valid sentence is found or the maximum number of iterations is reached (in this case, we accept the last generated sentence), as shown in Algorithm 1.
+Optionally, we may apply Principal Component Analysis (PCA) Jolliffe (2002) method to the embeddings to reduce their dimensionality before calculating the similarity [^3] In our experiments, this is the instruction in this case: “Represent the sentence for PCA:”.
 The reason for applying PCA is provided in Subsection 3.2.
 An overview of SimMark generation algorithm is depicted in the top part of Figure 2.
-The predefined interval is a hyperparameter chosen a priori by the user based on the distribution of similarities (or distances) between consecutive sentences’ embeddings generated by an unwatermarked LLM and human-written text.
+The predefined interval is a hyperparameter chosen a priori based on the distribution of similarities between consecutive sentences’ embeddings generated by an unwatermarked LLM and human-written text.
 The choice of this hyperparameter is critical for the performance of SimMark.
@@
 Furthermore, the choice of interval affects the robustness of SimMark against paraphrasing attacks.
-When an attacker paraphrases sentences, the similarities (or distances) may change and fall outside the interval.
-A larger interval provides greater robustness, as the watermark is less likely to be disrupted by paraphrasing.
+When an attacker paraphrases sentences, the similarities may change and fall outside the interval.
+Consequently, a larger interval provides greater robustness, as the watermark is less likely to be disrupted by paraphrasing.
 Therefore, the selection of the interval involves balancing several factors: it must not be too narrow to impede sentence generation while maintaining a low FP rate and adequate robustness against paraphrasing.
-Finding an optimal interval, or “sweet spot,” depends on the distribution of similarities between consecutive sentences, which varies across different embedding models.
-However, identifying such sweet spots is feasible when the user analyzes the similarity distributions of both LLM-generated and human-written text (See Appendix E for an example of finding the sweet spot).
+Finding a “sweet spot,” for the interval depends on the distribution of similarities between consecutive sentences, which can vary across models.
+However, identifying such sweet spots is feasible when we analyze the similarity distributions of both human-authored and LLM-generated text (see Appendix K for an example of finding a sweet spot).
 3.2 Watermarked Text Detection
-The detection of SimMark follows a similar methodology to KGW by employing a one-proportion $z$-test for hypothesis testing.
+The detection of SimMark follows a similar methodology to KGW by employing a $z$-test for hypothesis testing.
 However, akin to Hou et al.
-(2023, 2024), the detection operates at the sentence level rather than the token level.
-To perform detection, we first divide the input text into sentences [^2] Using sent_tokenize method of NLTK (Bird et al., 2009).
-and use the same Instructor-Large embedding model with identical instruction to compute the embeddings for each sentence.
-If PCA was used during the watermark generation process, it must also be applied here to ensure consistency.
-Algorithm 2 Detection Pseudo-Code for SimMark
-0:  input text, embedding model, predefined interval $[a,b]$, decay factor $K$, threshold $\beta$, optional PCA model
-1:  Split the input text into sentences excluding the first sentence (i.e, the prompt): $M=\{X_{2},\dots,X_{N_{\mathrm{total}}}\}$.
+(2024a), the detection operates at the sentence level rather than the token level.
+To perform detection, we first divide the input text into sentences and use the same semantic embedding model to compute the embeddings for each sentence.
+If PCA was applied during the watermarking process, it must also be applied during detection to ensure consistency.
+Algorithm 2 SimMark Detection Pseudo-Code
+0: input text, embedding model, interval $[a,b]$, decay factor $K$, threshold $\beta$
+1: Split the input text into sentences excluding the first sentence (i.e, the prompt): $M=\{X_{2},\dots,X_{N_{\mathrm{total}}}\}$, and set $N\leftarrow|M|$.
 2:for each sentence pair $(X_{i},X_{i+1})$ do
@@
 7:end for
-8:  $N\leftarrow|M|$
-9:  Soft count of valid sentences: $N_{\mathrm{valid\_soft}}\leftarrow\sum_{i=2}^{N+1}c_{i}$.
-10:  Estimate $p_{0}$ as the area under the human-written text embeddings similarity distribution curve within $[a,b]$.
-11:  Compute $z_{\mathrm{soft}}$ (i.e., the soft-$z$-score) using Eq. (2).
-12:  if $z_{\mathrm{soft}}>\beta$ then
-13:     Reject $H_{0}$, i.e., the input text is likely generated by an LLM watermarked by SimMark.  [Uncaptioned image]
-14:  else
-15:     Accept $H_{0}$, i.e., the input text is likely written by a human.   [Uncaptioned image]
-16:  end if
-Next, we compute the similarity (or distance) between consecutive sentences and count the number of valid sentences, denoted as $N_{\mathrm{valid\_soft}}$.
-A sentence $X_{i+1}$ is deemed valid if $s_{i+1}$ the similarity (or distance) between $X_{i+1}$ and $X_{i}$ lies within the predefined user interval $[a,b]$.
-However, paraphrasing attacks may alter embeddings significantly, causing the similarity (or distance) to deviate from the desired interval.
+8: Soft count of valid sentences: $N_{\mathrm{valid\_soft}}\leftarrow\sum_{i=2}^{N+1}c_{i}$.
+9: Estimate $p_{0}$ as the area under the human-written text embeddings similarity distribution curve within $[a,b]$.
+10: Compute $z_{\mathrm{soft}}$ using Eq. (2).
+11:if $z_{\mathrm{soft}}>\beta$ then
+12:  Reject $H_{0}$, i.e., text is likely generated by an LLM watermarked by SimMark.  [Uncaptioned image]
+13:else
+14:  Accept $H_{0}$, i.e., text is likely human-written.
+15:end if
+Next, we compute the similarity ($s_{i+1}=sim(e_{i},e_{i+1})$) between consecutive sentences ($X_{i}$ and $X_{i+1}$) and count the number of valid sentences ($N_{\mathrm{valid\_soft}}$).
+Sentence $X_{i+1}$ is deemed valid if $s_{i+1}$ lies within the predefined interval $[a,b]$.
+However, paraphrasing may alter embeddings significantly, causing the similarity to deviate from the desired interval.
 To mitigate this, we adopt a soft counting approach, where a sentence is considered partially valid if its similarity is near the interval boundaries.
 Specifically, the soft count of $X_{i+1}$, denoted as $c_{i+1}$, is defined as follows:
-$c_{i+1}=c(s_{i+1})=\begin{cases}1&\text{if }s_{i+1}\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\in[a,b]},\\ e^{-K\min\{|a-s_{i+1}|,|b-s_{i+1}|\}}&{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\text{otherwise.}}\end{cases}\end{aligned}$
-Here, $s_{i+1}=sim(e_{i},e_{i+1})$ represents the similarity (or distance) between consecutive sentence embeddings, and $K>0$ is a decay factor controlling the smoothness of the soft counting mechanism.
+$\begin{aligned} c_{i+1}=c(s_{i+1})=\begin{cases}1&\text{if }s_{i+1}\;{\color[rgb]{0,0.5,0}\in[a,b]},\\ e^{-K\min\{|a-s_{i+1}|,|b-s_{i+1}|\}}&{\color[rgb]{1,0,0}\text{otherwise.}}\end{cases}\end{aligned}$
+Here, $K>0$ is a decay factor controlling the smoothness of the soft counting.
 A higher $K$ makes the function behave closer to a step function, while a lower $K$ allows for smoother transitions, tolerating minor deviations outside the interval $[a,b]$.
-The total number of valid sentences is then computed as $N_{\mathrm{valid\_soft}}=\sum c_{i}$.
-Refer to Appendix C for experimental results demonstrating how this soft counting mechanism improves robustness against paraphrasing, with only a minimal impact on performance in non-paraphrased scenarios.
-Allowing for some degree of error through soft counting enables better detection under adversarial conditions while maintaining high accuracy in non-paraphrased cases.
-During our initial experiments, we observed that contrary to cosine similarity, Euclidean distance is very sensitive to paraphrasing and even a subtle change in the sentences would result in a huge difference in the distances of embeddings.
-We hypothesize that Euclidean distance is sensitive to noise in high-dimensional spaces such as the semantic space of the embedder.
-To mitigate this, we propose using PCA: We first train a PCA model on a dataset of human-written texts, to find the principal components of the sentence embeddings, i.e., the components that contribute the most to the semantical representation of the sentences.
-Then, we apply PCA to the embeddings to reduce their dimension.
-More details on the dimensionality reduction are provided in Section 4 and Appendix D.
-With this approach reverse-engineering the watermark algorithm becomes even harder as the third party who tries to reverse-engineer the algorithm, not only requires to know the embedder model and the instruction used for it, but it also needs to know the specifics of PCA setting (such as the number of principal components) and needs access to the dataset for training the PCA; otherwise, the similarity distribution of sentence embeddings would be different, making reverse engineering challenging.
-We found that this approach improved the robustness against paraphrasing when Euclidean distance is used.
-However, it would decrease the performance when cosine similarity is used.
-Returning to the detection, the null hypothesis $H_{0}$ is defined as follows:
-$H_{0}$: The sentences are written by humans, i.e., the text sequence is written (or generated) without knowledge of the valid interval in the similarity distribution of sentence embeddings.
+The total number of valid sentences is then computed as $N_{\mathrm{valid\_soft}}=\sum_{i}c_{i}$.
+Refer to Appendix I for an ablation study on how this approach improves robustness against paraphrasing, with only a minimal impact on performance in non-paraphrased scenarios, by allowing for some degree of error.
+During our initial experiments, we observed that, contrary to cosine similarity, Euclidean distance is very sensitive to paraphrasing, and even a subtle change in the sentences would result in a huge difference in the distances of embeddings.
+We hypothesize that Euclidean distance is sensitive to noise in high-dimensional spaces such as the semantic space of an embedder.
+To mitigate this, we propose using PCA: We fit a PCA model on a dataset of human-written texts to find the principal components of the sentence embeddings, i.e., the components that contribute the most to the semantical representation of the sentences.
+Then, we apply PCA to reduce embeddings’ dimension.
+More details on the dimensionality reduction are provided in Section 4 and Appendix J.
+This approach can make reverse-engineering more difficult, as it would require knowledge of not only the embedder model, but also the PCA setting (e.g., number of components, access to the dataset used for fitting it, etc.).
+Without all these details, reproducing the similarity distribution becomes less straightforward.
+The null hypothesis $H_{0}$ is defined as follows:
+$H_{0}$: The sentences are written by humans, i.e., the text sequence is generated without knowledge of the valid interval in the similarity of sentence embeddings.
 We calculate the $z$-statistic for the one-proportion $z$-test using the sample proportion $p=\frac{N_{\mathrm{valid\_soft}}}{N}$, where $N$ is the total number of samples (sentences).
-Since $N_{\mathrm{valid\_soft}}$ is a soft count of valid sentences, we refer to it as soft-$z$-score, which is given by $z_{\mathrm{soft}}=\frac{p-p_{0}}{\sqrt{\frac{p_{0}(1-p_{0})}{N}}}$.
-Here, $p_{0}$ is the population proportion and represents the ratio of valid sentences to all sentences in human-written text (i.e., in a text with no watermark), which is estimated as the area under the similarity distribution curve of consecutive human-written sentences within the interval $[a,b]$ (Like the one in Figure 6).
-Alternatively, the soft-$z$-score can be expressed as:
+Since $N_{\mathrm{valid\_soft}}$ is a soft count of valid sentences, we refer to it as soft-$z$-score, which is given by $z_{\mathrm{soft}}=\frac{p-p_{0}}{\sqrt{\frac{p_{0}(1-p_{0})}{N}}}$ or alternatively:
 $z_{\mathrm{soft}}=\tfrac{N_{\mathrm{valid\_soft}}-p_{0}N}{\sqrt{p_{0}(1-p_{0})N}}.$ (2)
-In this form, the value of $z_{\mathrm{soft}}$ can be interpreted as a normalized deviation of the number of valid sentences $N_{\mathrm{valid\_soft}}$ from its expectation $p_{0}N$.
+Here, the population proportion $p_{0}$ represents the ratio of valid sentences to all sentences in human-written text (i.e., a text with no watermark), which is estimated as the area under the similarity distribution curve of consecutive human-written sentences within the interval $[a,b]$ (like the one in Figure 9 in Appendix K).
+The value of $z_{\mathrm{soft}}$ can be interpreted as a normalized deviation of the number of valid sentences $N_{\mathrm{valid\_soft}}$ from its expectation $p_{0}N$.
 As highlighted in Algorithm 2, the null hypothesis $H_{0}$ is rejected if $z_{soft}>\beta$, where $\beta$ is a threshold determined empirically by running the detection algorithm on human-written text.
 The threshold $\beta$ is selected to maintain a desired FP rate (i.e., minimizing the misclassification of human-written text as LLM-generated).
-Details on the computation of $\beta$ are provided in the Appendix F.
-4 Experiments
+Details on the computation of $\beta$ are provided in Appendix L.
+4 Experiments & Results
 Dataset Algorithm No Paraphrase Pegasus Pegasus-Bigram Parrot Parrot-Bigram GPT3.5 GPT3.5-bigram Avg. Paraphrased
-RealNews KGW (Kirchenbauer et al.) 99.6 / 98.4 / 98.9 95.9 / 82.1 / 91.0 92.1 / 42.7 / 72.9 88.5 / 31.5 / 55.4 83.0 / 15.0 / 39.9 82.8 / 17.4 / 46.7 75.1 / 5.9 / 26.3 86.2 / 32.4 / 55.4
+RealNews
+UW (Zhao et al.) 99.9 / 99.1 / 99.9 98.5 / 85.6 / 95.3 97.9 / 73.5 / 91.7 97.9 / 70.9 / 91.9 97.4 / 62.8 / 89.4 97.4 / 59.1 / 87.9 93.7 / 37.0 / 70.8 97.1 / 64.8 / 87.8
+KGW (Kirchenbauer et al.) 99.6 / 98.4 / 98.9 95.9 / 82.1 / 91.0 92.1 / 42.7 / 72.9 88.5 / 31.5 / 55.4 83.0 / 15.0 / 39.9 82.8 / 17.4 / 46.7 75.1 / 5.9 / 26.3 86.2 / 32.4 / 55.4
 SIR (Liu et al.) 99.9 / 99.4 / 99.9 94.4 / 79.2 / 85.4 94.1 / 72.6 / 82.6 93.2 / 62.8 / 75.9 95.2 / 66.4 / 80.2 80.2 / 24.7 / 42.7 77.7 / 20.9 / 36.4 89.1 / 54.4 / 67.2
@@
 Euclidean-SimMark [^7] PCA is applied. (ours) 99.8 / 98.5 / 99.3 97.2 / 72.3 / 89.1 96.9 / 70.0 / 87.4 95.7 / 60.2 / 82.5 95.7 / 59.1 / 81.5 94.1 / 51.6 / 76.2 88.2 / 29.7 / 53.5 94.6 / 57.2 / 78.4
-BookSum KGW 99.6 / 99.0 / 99.2 97.3 / 89.7 / 95.3 96.5 / 56.6 / 85.3 94.6 / 42.0 / 75.8 93.1 / 37.4 / 71.2 87.6 / 17.2 / 52.1 77.1 / 4.4 / 27.1 91.0 / 41.2 / 67.8
+BookSum
+UW 100 / 100 / 100 99.5 / 89.8 / 98.5 98.6 / 71.2 / 93.0 98.9 / 79.4 / 94.8 98.6 / 72.1 / 92.9 93.2 / 24.6 / 57.9 86.0 / 9.2 / 30.5 95.8 / 57.7 / 77.9
+KGW 99.6 / 99.0 / 99.2 97.3 / 89.7 / 95.3 96.5 / 56.6 / 85.3 94.6 / 42.0 / 75.8 93.1 / 37.4 / 71.2 87.6 / 17.2 / 52.1 77.1 / 4.4 / 27.1 91.0 / 41.2 / 67.8
 SIR 100 / 99.8 / 100 93.1 / 79.3 / 85.9 93.7 / 69.9 / 81.5 96.5 / 72.9 / 85.1 97.2 / 76.5 / 88.0 80.9 / 39.9 / 23.6 75.8 / 19.9 / 35.4 89.5 / 59.7 / 66.6
@@
 Euclidean-SimMark (ours) 100 / 100 / 100 98.8 / 82.6 / 94.9 98.6 / 80.4 / 93.4 97.9 / 75.3 / 91.1 97.9 / 73.3 / 91.6 99.7 / 94.4 / 98.8 99.5 / 91.9 / 97.6 98.7 / 83.0 / 94.6
-Reddit-TIFU KGW 99.3 / 97.5 / 98.1 94.1 / 87.2 / 87.2 91.7 / 67.2 / 67.6 79.5 / 22.8 / 43.3 82.8 / 27.6 / 49.7 84.1 / 27.3 / 50.9 79.8 / 19.3 / 41.3 85.3 / 41.9 / 56.7
+Reddit-TIFU
+UW 99.9 / 99.5 / 99.8 97.3 / 73.4 / 91.1 94.1 / 48.3 / 77.2 90.6 / 37.1 / 64.0 89.2 / 33.7 / 60.4 86.3 / 26.9 / 52.9 74.3 / 13.2 / 30.0 88.6 / 38.8 / 62.6
+KGW 99.3 / 97.5 / 98.1 94.1 / 87.2 / 87.2 91.7 / 67.2 / 67.6 79.5 / 22.8 / 43.3 82.8 / 27.6 / 49.7 84.1 / 27.3 / 50.9 79.8 / 19.3 / 41.3 85.3 / 41.9 / 56.7
+SIR 99.6 / 97.2 / 99.7 90.0 / 48.7 / 77.4 90.9 / 33.1 / 71.1 87.1 / 15.0 / 50.9 86.9 / 12.8 / 49.8 91.1 / 15.0 / 61.4 84.3 / 5.5 / 39.1 88.4 / 21.7 / 58.3
 SemStamp 99.7 / 97.7 / 98.2 98.4 / 92.8 / 95.4 98.0 / 89.0 / 92.9 90.2 / 56.2 / 70.5 93.9 / 71.8 / 82.3 87.7 / 47.5 / 58.2 87.4 / 43.8 / 55.9 92.6 / 66.9 / 75.9
@@
 Euclidean-SimMark [^7] PCA is applied. (ours) 99.8 / 98.7 / 99.2 99.0 / 94.7 / 97.6 99.0 / 91.9 / 96.2 97.8 / 75.9 / 89.5 97.7 / 76.4 / 90.4 98.7 / 83.7 / 95.2 96.8 / 65.8 / 87.3 98.2 / 81.4 / 92.7
-[…]
-[…]
-No watermark 11.89 11.43 3.10
-SemStamp 12.89 11.51 3.17
-$k$-SemStamp 11.82 11.48 3.11
-Cosine-SimMark (Ours) 12.69 11.50 3.39
-Euclidean-SimMark [^7] PCA is applied. (Ours) 9.67 11.50 3.28
-Table 2: Comparison of the quality of text generated using different watermarking algorithms on the BookSum dataset.
-For PPL metric, lower values indicate better quality, while higher values are better for Ent-3 and Sem-Ent.
-SimMark watermarking achieves quality comparable to no watermarking, demonstrating minimal impact on text quality.
-In this section, we evaluate the performance of our watermarking algorithm, SimMark, across different datasets.
-Performance is assessed using the area under the receiver operating characteristic curve (ROC-AUC) and true positive rate (TP) at fixed false positive rates (FP) of 1% and 5%, (TP@1%FP and TP@5%FP).
-Higher values indicate better performance across all metrics.
-For dimensionality reduction, we trained a PCA model on 8000 samples from the RealNews subset of the C4 dataset (Raffel et al., 2020), reducing embedding dimensions from 768 to 16.
+Table 1: Performance of different algorithms across datasets and paraphrasers, evaluated using ROC-AUC $\uparrow$ / TP@FP=1% $\uparrow$ / TP@FP=5% $\uparrow$, respectively (↑: higher is better).
+In each column, bold value indicates the best performance for a given dataset and metric, while underlined value denotes the second-best.
+SimMark consistently outperforms or is on par with other state-of-the-art methods across datasets, paraphrasers, and is the best on average.
+Algorithm PPL $\downarrow$ Ent-3 $\uparrow$ Sem-Ent $\uparrow$
+No watermark 11.89 11.43 3.32
+UW 14.57 11.47 3.33
+KGW 14.92 11.32 2.95
+SIR 20.34 11.57 3.18
+SemStamp 12.89 11.50 3.32
+$k$-SemStamp 11.82 11.48 3.32
+SimMark (ours) 12.69 11.50 3.37
+Table 2: Comparison of the quality of text watermarked using different algorithms on the BookSum dataset (↓: lower is better, ↑: higher is better).
+SimMark yields quality metrics comparable to the no-watermark baseline, indicating minimal impact on text quality and semantic diversity.
+In contrast, token-level methods (UW, KGW, and SIR) notably degrade the text quality, especially in terms of perplexity.
+Performance of SimMark is evaluated across different datasets and models using the area under the receiver operating characteristic curve (ROC-AUC) and true positive rate (TP) at fixed FP rates of 1% and 5% (TP@1%FP and TP@5%FP).
+Higher values indicate better performance across all metrics [^4] Like Hou et al.
+(2024a), all results are from a single run..
+For dimensionality reduction, we fitted a PCA model on 8000 samples from the RealNews subset of the C4 dataset Raffel et al.
+(2020), reducing embedding dimensions from 768 to 16.
 After testing various principal component counts (ranging from 512 to 16), we found 16 to yield the best results.
-PCA improved robustness against paraphrasing attacks when Euclidean distance was used (except on the BookSum dataset) but consistently degraded performance when cosine similarity was employed across all datasets.
-During our experiments, we applied PCA wherever it enhanced performance.
-More details and the results of these experiments are summarized in Table 4 in Appendix D.
-Across all experiments, the decay factor was set to $K=250$, as this value provided an optimal trade-off between performance under both non-paraphrased and paraphrased conditions (See Appendix C).
-The threshold $\beta$ was determined programmatically during the detection (See Appendix F) to achieve the specified FP rates (1% or 5%).
-The predefined intervals $[0.68,0.76]$ for cosine similarity and $[0.28,0.36]$ for Euclidean distance with PCA and $[0.4,0.55]$ for Euclidean distance without PCA were found to be near-optimal.
-While the intervals could have been further optimized for each dataset individually, we chose not to do so to demonstrate the general performance of our method.
-For additional details on the experimental environment refer to Appendix H.
+During our experiments, we evaluated both settings (with and without PCA).
+Specifically, PCA improved robustness against paraphrasing attacks when Euclidean distance was used (except on the BookSum dataset), but consistently degraded performance when cosine similarity was employed across all datasets.
+The results of these experiments are summarized in Table 7 in Appendix J.
+Across all experiments, the decay factor was set to $K=250$, as this value provided an optimal trade-off between performance under both non-paraphrased and paraphrased conditions (see Appendix I for an ablation study on this).
+The threshold $\beta$ was determined empirically during the detection (refer to Appendix L for details) to achieve the specified FP rates (1% or 5%).
+The intervals $[0.68,0.76]$ for cosine similarity and $[0.28,0.36]$ for Euclidean distance with PCA and $[0.4,0.55]$ for Euclidean distance without PCA were found to be near-optimal, as detailed in Appendix K.
+While the intervals as well as other hyperparameters such as decay factor $K$ could have been further optimized for each dataset or paraphraser individually, we chose not to do so to show the general performance of our method (in contrast, $k$-SemStamp fine-tunes domain-specific embedders that are optimized per setting).
 4.1 Models and Datasets
-To ensure comparability with prior works by Hou et al.
-(2023, 2024), we also employed the same fine-tuned version of OPT-1.3B (Zhang et al., 2022), that they used as our LLM [^3] Used AbeHou/opt-1.3b-semstamp model..
-For semantic embedding, we utilized the Instructor-Large model [^4] Used hkunlp/instructor-large embedding model.
-(Su et al., 2023), as mentioned earlier.
-Following Kirchenbauer et al.
-(2023), sampling from the LLM was performed with a temperature of 0.7 and a repetition penalty of 1.05, while the minimum and the maximum number of generated tokens were set to 195 and 205, respectively.
-The maximum number of iterations for rejection sampling is set to 100, following Hou et al.
-(2023, 2024).
-In our experiments, we used the RealNews subset of the C4, the BookSum (Kryscinski et al., 2022), and Reddit-TIFU (Kim et al., 2019) datasets, as in Hou et al.
+For our experiments, we used the same fine-tuned version of OPT-1.3B Zhang et al.
+(2022) as in Hou et al.
+(2024a, b) [^5] Used AbeHou/opt-1.3b-semstamp (1.3B) model.
+to ensure fair comparison.
+However, we emphasize that our method is model-agnostic and it treats the LLM as a black-box text generator.
+As such, if the method performs well on one family of models, it is expected to generalize to others.
+To support this, we also tested our method on Gemma3-4B model Team et al.
+(2025) and observed similar results (see Appendix D).
+For semantic embedding, we utilized Instructor-Large [^6] Used hkunlp/instructor-large (335M) model.
+Su et al.
 (2023).
-1000 samples from each dataset were chosen to analyze the detection results and the text quality.
-Each sample was segmented into sentences, with the first sentence serving as the prompt to the LLM.
+Appendix B includes additional details on the experimental configurations.
+In our experiments, we used three English datasets: RealNews subset of C4777Dataset card: allenai/c4 (Validation split) Raffel et al.
+(2020), BookSum [^8] Dataset card: kmfoda/booksum (Validation split) Kryscinski et al.
+(2022), and Reddit-TIFU [^9] Dataset card: ctr4si/reddit_tifu (Train split, short subset) Kim et al.
+(2019) datasets, as in Hou et al.
+(2024a).
+Specifically, 1000 samples from each dataset were chosen to analyze the detection performance and the text quality.
+Each sample was segmented into sentences [^10] Using sent_tokenize method of NLTK Bird et al.
+(2009)., with the first sentence serving as the prompt to the LLM.
 We evaluated text quality after applying SimMark using the following metrics:
-  • Perplexity (PPL): Measures how surprising the text is to a larger LLM (OPT-2.7B (Zhang et al., 2022)). Lower values indicate higher quality.
-  • Tri-gram Entropy (Ent-3) (Zhang et al., 2018): To assess textual diversity via the entropy of the tri-grams distribution.
-  • Semantic Entropy (Sem-Ent) (Han et al., 2022): Used to measure semantic informativeness and diversity of the text.
-4.2 Paraphrasing Attack
-To evaluate the robustness of SimMark, we tested it against paraphrasing attacks using three models: Pegasus paraphraser [^5] Used tuner007/pegasus_paraphrase paraphrasing model.
-(Zhang et al., 2020), Parrot paraphraser (Damodaran, 2021), and GPT-3.5-Turbo (OpenAI, 2022).
+  • Tri-gram Entropy (Ent-3) ↑ Zhang et al. (2018): Assesses textual diversity via the entropy of the tri-grams distribution.
+  • Semantic Entropy (Sem-Ent) ↑ Han et al. (2022): Measures semantic informativeness and diversity of the text.
+  • Perplexity (PPL) ↓ Jelinek et al. (1977): Measures how surprising the text is to an oracle LLM [^11] Used facebook/opt-2.7b, following Hou et al. (2024a, b)..
+4.2 Paraphrase Attack
+To evaluate the robustness of SimMark against paraphrase attacks, we tested it using three paraphrasers: $I.$ Pegasus paraphraser [^12] Used tuner007/pegasus_paraphrase (568M) model.
+Zhang et al.
+(2020), II.II.
+Parrot paraphraser [^13] Used parrot_paraphraser_on_T5 (220M) model.
+Damodaran (2021), $III.$ GPT-3.5-Turbo OpenAI (2022).
+Refer to Appendix N to find the prompts used with GPT-3.5-Turbo.
 Kirchenbauer et al.
 (2024) observed that prompting models to paraphrase entire texts often results in summarized outputs, with the summarization ratio worsening for longer inputs.
-To prevent such summarization, we adopted a sentence-by-sentence paraphrasing scheme, which also ensures our results are comparable to prior works such as Hou et al.
-(2023, 2024).
-The quality of paraphrases was assessed using BertScore [^6] Used microsoft/deberta-xlarge-mnli (He et al., 2021) model.
-(Zhang* et al., 2020), with all settings consistent with prior works of Hou et al.
-(2023, 2024).
+To prevent any information loss caused by such summarization, we adopted a sentence-by-sentence paraphrasing scheme, which also ensures our results are comparable to Hou et al.
+(2024a, b).
+The quality of paraphrases was assessed using BertScore [^14] Used deberta-xlarge-mnli (750M) He et al.
+(2021).
+Zhang* et al.
+(2020), with all settings consistent with Hou et al.
+(2024a, b).
 The bottom part of Figure 2 demonstrates the paraphrase attack and detection phase in more detail.
-Refer to Appendix I to find the prompts used with GPT-3.5-Turbo.
-We also included the results for the bigram paraphrase attacks introduced by Hou et al.
-(2023), with identical settings (25 rephrases for each sentence, etc.).
-This type of attack involves generating multiple paraphrases for each sentence, and choosing the one that increases the likelihood of disrupting statistical signatures embedded in the text, especially for token-level watermarking algorithms (Hou et al., 2023).
-While this type of attack significantly impacts most other models, SimMark demonstrates greater robustness against it.
-We must highlight that direct comparisons to $k$-SemStamp are not entirely fair, as $k$-SemStamp relies on domain-specific clustering of semantic spaces, making it domain-dependent.
-In contrast, SimMark, similar to SemStamp, is domain-independent.
-Still, SimMark outperforms $k$-SemStamp in most cases across various datasets and metrics, further underscoring its universality and robustness.
-4.3 Paraphrase+Drop Attack
-Figure 3: Detection results (ROC-AUC) under the Paraphrase+Drop Attack.
-Pegasus is used as the paraphraser.
-Here, 0 probability means no sentence is dropped, i.e., sentences are only paraphrased.
-SimMark is more robust than SemStamp and $k$-SemStamp across varying sentence drop probabilities.
-To further evaluate the robustness of our method and its real-world applicability, we introduce a more challenging attack scenario, termed the Paraphrase+Drop Attack.
-We argue that an adversary not only paraphrases the text but also drops sentences deemed redundant.
-This reflects common editing practices since LLMs are known for their tendency to generate lengthy outputs with redundancy, likely due to their imperfect reward modelling (Chiang & Lee, 2024).
-To simulate this scenario, we first paraphrase the text using the method described in the previous subsection and then randomly drop sentences from the paraphrased text.
-This combined attack tests the resilience of SimMark under more aggressive transformations that mimic real-world adversarial editing.
+We also included the results for the bigram paraphrase attack introduced by Hou et al.
+(2024a), with identical settings (25 rephrases for each sentence when using Pegasus and Parrot, etc.).
+This attack involves generating multiple paraphrases for each sentence, and choosing the one that increases the likelihood of disrupting statistical signatures embedded in the text, especially for token-level algorithms Hou et al.
+(2024a).
+While this attack significantly impacts most other methods, SimMark demonstrates greater robustness against it.
+We must highlight that $k$-SemStamp relies on domain-specific clustering of semantic spaces, making it domain-dependent.
+In contrast, both SimMark and SemStamp are domain-independent.
+Still, SimMark outperforms both in nearly all cases across various datasets and metrics, further underscoring its universality and robustness.
+4.3 Robustness to Sentence-Level Perturbations
+To further evaluate the robustness of our method and its real-world applicability, we introduce more challenging attack scenarios: Paraphrase+Drop and Paraphrase+Merge Attacks.
+These scenarios simulate realistic adversarial editing strategies where a user attempts to remove or merge sentences after paraphrasing, while maintaining fluency.
+Paraphrase+Drop assumes that an adversary not only paraphrases the text but also drops sentences deemed redundant.
+This reflects common editing practices, especially since LLMs often produce verbose outputs due to imperfect reward modeling Chiang and Lee (2024).
+To simulate this, we first paraphrase the input text and then randomly drop sentences with a specified probability $p$.
+Similarly, Paraphrase+Merge is designed to test robustness under more subtle structural changes.
+After paraphrasing, we replace end-of-sentence punctuations (e.g., ., ?, !) with the word “and” with probability $0<p<\frac{1}{2}$.
+In our experiments, we avoid higher values of $p$ as they result in unnaturally long and less fluent sentences.
+This setup simulates a realistic scenario where an adversary attempts to merge sentences while preserving the overall coherence of the text.
+These combined attacks serve to stress-test the resilience of watermarking methods under more naturalistic adversarial conditions that go beyond simple paraphrasing.
 4.4 Results & Discussion
-We compared the performance of SimMark against state-of-the-art watermarking algorithms through extensive experiments.
+Figure 3: Detection performance of different watermarking methods under various paraphrasing attacks, measured by TP@FP=1% ↑ and averaged across all three datasets (RealNews, BookSum, Reddit-TIFU).
+Each axis corresponds to a specific paraphrasing attack method (e.g., Pegasus-Bigram), and higher values are better.
+Our methods, cosine-SimMark and Euclidean-SimMark, consistently outperform or match baselines across most paraphrasers, especially under more challenging conditions such as bigram-level paraphrasing.
+We compared the performance of SimMark against SOTA watermarking algorithms through extensive experiments.
 Our primary baseline was SemStamp, a sentence-level semantic watermarking method.
-We also included $k$-SemStamp, an improvement over SemStamp tailored to specific text domains.
-Results for SemStamp, $k$-SemStamp, and other methods, including KGW and SIR, were extracted directly from Hou et al.
-(2024).
-Table 1 presents detection performance before and after paraphrasing attacks, while Table 2 provides text quality evaluation results.
-Our algorithm (the last two rows in each section) does not affect the text quality while being effective and consistently outperforming or matching other state-of-the-art methods, achieving the highest average performance across all paraphrasers and datasets.
-Notably, domain-independent SimMark surpasses our primary baseline, SemStamp, and matches or exceeds the domain-dependent $k$-SemStamp.
-Refer to Appendix B for qualitative examples of SimMark.
-A key aspect to consider is that the fine-tuning of SemStamp and $k$-SemStamp’s embedding model on text paraphrased by Pegasus likely contributes to its improved robustness against this specific paraphraser but may introduce bias.
-Additionally, the results for the Reddit-TIFU dataset were only available for SemStamp and not $k$-SemStamp, likely due to the dataset’s informal diverse text style and $k$-SemStamp’s limitation for text to belong to a specific domain such as news articles or scientific writings (Hou et al., 2024).
-Figure 3 presents the detection performance (in terms of ROC-AUC) of SimMark, $k$-SemStamp, and SemStamp against Paraphrase+Delete Attack under varying sentence drop probabilities, evaluated on the RealNews and BookSum datasets.
-The results demonstrate that SimMark consistently achieves higher performance compared to other methods, even as the sentence drop probability increases.
-These findings underscore the robustness of SimMark in handling more challenging real-world adversarial scenarios.
-Despite Hou et al.
-(2023, 2024) releasing their code and data, we were unable to reproduce their reported results fully.
-Consequently, there are minor discrepancies between our reproduction results (shown in Figure 3 for the case with zero probability) and those presented in Table 1, which were extracted directly from their paper.
-Notably, the results for $k$-SemStamp on the BookSum dataset differed substantially, leading us to exclude them from Figure 3.
-Additionally, the results reported in Table 2 (our reproduction) also show slight differences from their paper, likely due to hyperparameter details that were not explicitly documented.
-Regarding sampling efficiency, for BookSum dataset for instance, cosine-SimMark required an average of 7.1 samples per sentence from the LLM, compared to $k$-SemStamp and SemStamp, which averaged 13.3 and 20.9 samples, respectively (Hou et al., 2024).
-This demonstrates that our method not only outperforms these baselines but is also significantly (about 2-3 times) more efficient.
-Further analysis of this, including theoretical estimates is provided in Appendix G.
+We also included $k$-SemStamp, an improvement over SemStamp tailored to specific domains.
+Results for SemStamp, $k$-SemStamp, KGW, and SIR, were extracted directly from Hou et al.
+(2024a, b) [^15] Despite Hou et al.
+(2024a, b) releasing their code and data, we were unable to reproduce their reported results fully.
+Consequently, there are minor discrepancies between our reproduction results (shown in Figure 4 for cases with $p=0$) and those presented in Table 1 (extracted directly from their paper).
+Additionally, the results reported in Table 2 (our reproduction) also show slight differences from their paper, likely due to hyperparameter details that were not explicitly documented..
+Table 1 presents detection performance pre and post paraphrase attacks, while Table 2 provides text quality evaluation results.
+To better visualize robustness across paraphrasing attacks, we aggregate TP@FP=1% scores from Table 1 across datasets and present them in a radar plot (Figure 3).
+Additional radar plots and a summary table comparing performance across all paraphrasing settings are provided in Appendix C (see Table 3 and Figure 5).
+We also conducted a small-scale A/B test to assess watermark imperceptibility.
+To see the results, please refer to Appendix M.
+Overall, our algorithm was imperceptible to the evaluators, and impacted the text quality minimally while being effective and consistently outperforming or matching other SOTA methods, achieving the highest average paraphrased performance across all datasets.
+Notably, our method, SimMark (domain-independent), surpasses the primary baseline, SemStamp (domain-independent), and is on par with or exceeds $k$-SemStamp (domain-dependent).
+A key aspect to consider is that the fine-tuning of SemStamp and $k$-SemStamp’s embedding model on text paraphrased by Pegasus likely contributes to their higher robustness against Pegasus but may introduce bias and reduce general applicability (as shown in Table 2 of Hou et al.
+(2024b), $k$-SemStamp loses performance under domain shift).
+Additionally, the results for the Reddit-TIFU dataset were only available for SemStamp and not $k$-SemStamp, likely due to the dataset’s informal, diverse text style and $k$-SemStamp’s limitation for text to belong to a specific domain, such as news articles or scientific writings Hou et al.
+(2024b).
+While SimMark demonstrates strong performance, we acknowledge that it is not always the best across every setting.
+SimMark is designed with generality, robustness, and black-box compatibility in mind.
+It forgoes domain-specific fine-tuning in favor of broader applicability, which may explain why it does not always outperform more narrowly optimized methods.
+Several factors may explain its relative underperformance in some cases:
+First, some variability arises from randomness in datasets and generation.
+LLMs rely on pseudo-random decoding, which may introduce subtle fluctuations in outputs.
+More importantly, as SimMark operates at the sentence level, it cannot fine-tune token-level patterns similar to token-level methods (e.g., UW), which may subtly manipulate every token to embed statistical signals.
+This granularity difference can be particularly advantageous when generation length is short (e.g., a 200-token cap).
+In such cases, token-level methods benefit from having more embedding capacities, whereas sentence-level methods like SimMark may yield weaker statistical signals, resulting in greater performance variability in shorter texts.
+In other words, token-level methods benefit from high-frequency watermark signals, whereas SimMark injects at a coarser granularity.
+As a result, for SimMark, detection performance improves with longer generations, where more sentence-level signal can accumulate.
+Figure 4: Detection performance (ROC-AUC ↑) under two adversarial settings using RealNews dataset.
+Left: Paraphrase+Drop Attack, where random sentences are removed after paraphrasing.
+SimMark, in nearly all parameter regimes, outperforms other methods under this attack.
+Right: Paraphrase+Merge Attack, where sentence boundaries (punctuations) are probabilistically replaced with “and” to merge sentences.
+Although SimMark performs best among sentence-level approaches, UW remains highly robust due to its token-level nature.
+Figure 4 presents the ROC-AUC performance of UW (the best token-level method in our experiments), SimMark, $k$-SemStamp, and SemStamp under Paraphrase+Drop and Paraphrase+Merge attacks, evaluated on the RealNews dataset.
+Under Paraphrase+Drop, across most parameter regimes, SimMark outperforms all other methods, sustaining higher detection performance even as the attack intensity increases.
+While SimMark shows superior performance among sentence-based methods under Paraphrase+Merge, UW maintains the highest robustness because it operates at the token level.
+Regarding sampling efficiency, for BookSum dataset for instance, SimMark required an average of 7.1 samples per sentence from the LLM, compared to $k$-SemStamp and SemStamp, which averaged 13.3 and 20.9 samples, respectively Hou et al.
+(2024b).
+This demonstrates that our method not only outperforms these baselines but is also 2-3 times more efficient (see Appendix G for theoretical estimates of this).
+For a comparison between SimMark and token-level methods in terms of real-world runtime, see Appendix F.
+Finally, refer to Appendix H for qualitative examples of SimMark.
 5 Conclusion
-In this paper, we introduced SimMark, a similarity-based, sentence-level posthoc watermarking algorithm.
-Unlike existing approaches, SimMark operates without requiring access to the internal logits or architecture of the model, ensuring compatibility with a wide range of LLMs, including API-only models.
-By utilizing a general-purpose sentence embedding model and integrating a soft counting mechanism, SimMark combines robustness against paraphrasing with applicability to diverse domains.
-Furthermore, its resistance to reverse engineering ensures resilience against spoofing attacks.
-Experimental results show that SimMark outperforms state-of-the-art watermarking algorithms, such as SemStamp, in both efficiency and robustness to adversarial paraphrasing, representing a significant step forward in practical watermarking techniques for LLM-generated content.
-Due to space constraints, the discussion of the future directions is deferred to Appendix A.
-Acknowledgements
-This work was supported by the NSERC Discovery Grant No. RGPIN-2019-05448.
-Impact Statement [^7] Generative AI tools, such as ChatGPT (OpenAI, 2022), were used to refine this manuscript.
-The authors retain full responsibility for all content presented in the paper, ensuring adherence to academic integrity and ethical research practices.
-By enabling robust detection of LLM-generated text, particularly under paraphrasing attacks, SimMark addresses ethical concerns surrounding the transparency and accountability of AI-generated content.
-However, like any watermarking method, there are potential risks, such as falsely implicating human authors or adversaries developing more advanced techniques for spoofing attacks or bypassing detection.
+In this paper, we introduced SimMark, a similarity-based, robust sentence-level watermarking algorithm.
+Unlike existing approaches, SimMark operates without requiring access to the internals of the model, ensuring compatibility with a wide range of LLMs, including API-only models.
+By utilizing a pre-trained general-purpose embedding model and integrating a soft counting mechanism, SimMark combines robustness against paraphrasing with applicability to diverse domains.
+Experimental results show that SimMark outperforms SOTA sentence-level watermarking algorithms in both efficiency and robustness to paraphrasing, representing a step forward in fully semantic watermarking for LLMs.
+Limitations
+While SimMark demonstrates outstanding performance, there are still some areas that warrant further exploration:
+Rejection Sampling Overhead.
+The rejection sampling process requires generating multiple candidate sentences until a valid sentence is accepted.
+Although our method is significantly (2-3 times) more efficient than prior approaches such as SemStamp and $k$-SemStamp, there is still a notable decrease in generation speed due to rejection sampling.
+Techniques like batch sampling or parallel sampling could potentially mitigate this issue, though at the expense of higher computational resource usage.
+Future research should focus on optimizing the method to balance efficiency and resource requirements.
+Resistance to More Advanced Attacks.
+While SimMark demonstrates robustness against paraphrasing attacks, it may not be immune to more sophisticated adversarial transformations.
+In particular, detection could become less effective when watermarked text is interleaved with or embedded within a larger body of unwatermarked content.
+Additionally, although reverse engineering the exact watermarking rules is non-trivial, an adversary may attempt a spoofing attack by approximating our setup—for instance, by employing a publicly available embedding model or fitting a PCA model with publicly available datasets.
+While such attempts may not perfectly replicate the original embedding distribution, they could still pose a threat.
+We leave a thorough investigation into vulnerabilities and corresponding defences against reverse engineering to future work.
+Dependency on Predefined Intervals.
+In our experiments, we used consistent, predefined intervals across all datasets and observed consistently strong performance.
+Notably, we did not observe any noticeable degradation in text quality due to this interval constraint during rejection sampling (as shown in Table 2), likely because the constraint applies only to consecutive sentences.
+Nonetheless, slight variations in the embeddings similarity distribution of LLM-generated text across different models/datasets may impact watermarking effectiveness.
+Adaptive strategies for setting these intervals dynamically (or pseudo-randomly) could not only improve performance but also make reverse-engineering the algorithm more difficult.
+Ethical Considerations [^16] Generative AI tools, such as ChatGPT, were used to refine this manuscript.
+The authors retain full responsibility for all content presented in the paper, ensuring adherence to academic integrity and ethical research standards.
+Potential Risks.
+By enabling robust detection of LLM-generated text, particularly under paraphrasing attacks, SimMark tries to address ethical concerns surrounding the transparency and accountability of AI-generated content.
+However, like any watermarking algorithm, there are potential risks, such as falsely implicating human authors or adversaries developing more advanced techniques for spoofing attacks or bypassing detection.
 We acknowledge these limitations and advocate for the responsible deployment of such tools in combination with other verification mechanisms to mitigate these risks and ensure ethical, fair deployment.
@@
 (2021).
-Appendix A Future Work
-While SimMark demonstrates outstanding performance, there are still some areas that warrant further exploration:
-Reduce Rejection Sampling Overhead.
-The rejection sampling process requires generating multiple candidate sentences until a valid sentence is accepted.
-Although our method is 2-3 times more efficient than prior approaches such as SemStamp and $k$-SemStamp, there is still a notable decrease in generation speed due to rejection sampling.
-Techniques like batch sampling or parallel sampling could potentially mitigate this issue, though at the expense of higher computational resource usage.
-Future research should focus on optimizing these methods to balance efficiency and resource requirements.
-Improve Resistance to More Advanced Attacks.
-Although SimMark is robust against paraphrasing attacks to some extent, more advanced text transformations (e.g., sentence splitting and merging, semantic reordering) may still pose challenges.
-Additionally, our detection algorithm may struggle when watermarked text is mixed into a larger body of human-written content.
-Further research is needed to enhance resilience against a broader range of adversarial transformations.
-Setting the Interval Dynamically.
-In our experiments, we used consistent predefined intervals across all datasets and observed consistently strong performance.
-Notably, we did not observe any degradation in text quality due to this interval constraint during rejection sampling (as shown in Table 2), likely because the constraint applies only to consecutive sentences.
-Nonetheless, slight variations in the embeddings similarity distribution of LLM-generated text across different models and domains may impact watermarking effectiveness.
-Adaptive strategies for setting these intervals dynamically (or pseudo-randomly) could not only improve performance but also make reverse-engineering the algorithm even more difficult.
-Appendix B Examples of Watermarked Text
-Figures 4 and 5 provide examples of text generated with and without the SimMark watermark.
+Use of Models and Datasets.
+Our research used datasets and pretrained models from the Hugging Face Hub [^17] https://huggingface.co, a public platform hosting machine learning resources under various licenses.
+We adhered to all license terms and intended usage guidelines for each artifact, which are documented on their individual Hugging Face model or dataset cards cited in the paper.
+All resources were used solely for research purposes in accordance with their intended use and respective licenses.
+The datasets employed are publicly available, widely used in prior research, and, to the best of our knowledge, free of personally identifiable information or offensive content.
+Any outputs generated by our method are intended for academic use, not for real-world or commercial applications, and no personal or sensitive data was processed.
+Any new artifacts we create (e.g., watermarked samples) are intended solely for academic evaluation, and we do not release any derivative data that violates original licensing terms.
+Acknowledgments
+This work was supported by the NSERC Discovery Grant No. RGPIN-2019-05448.
+Supplemental Materials
+Appendix A Aditional Related Work
+A.1 Token-Level Watermarking
+Zhao et al.’s (2023) UNIGRAM-WATERMARK (UW) builds upon KGW by fixing the red and green lists instead of pseudo-randomly selecting them, proving that, compared to KGW, their method is more robust to paraphrasing and editing Zhao et al.
+(2023).
+However, as outlined by Hou et al.
+(2024a), this algorithm can be reverse-engineered, rendering it impractical for high-stakes, real-world applications (for details about the reverse-engineering procedure, refer to Hou et al.
+(2024a)).
+The Semantic Invariant Robust (SIR) watermark in Liu et al.
+(2023) is also similar to KGW but is designed to be less sensitive to attacks involving synonym replacement or advanced paraphrasing.
+SIR achieves this by altering the LLM logits based on the semantics of previously generated tokens, using a semantic embedding model to compute semantic representations, and training a model that adjusts LLM’s logits based on the semantic embeddings of prior tokens Liu et al.
+(2023).
+A.2 Post-Hoc Watermarking
+Chang et al.’s (2024) PostMark is a post-hoc watermarking algorithm designed to work without access to model logits, making it compatible with API-only LLMs.
+It constructs an input-dependent set of candidate words using semantic embeddings and then prompts another LLM (e.g., GPT-4o) to insert these words into the generated text.
+Detection relies on statistical analysis of the inserted words.
+While PostMark’s compatibility with black-box LLMs is a strength, the approach is expensive—watermarking 100 tokens is estimated to cost around \$1.2 USD Chang et al.
+(2024).
+Yang et al.
+(2023) propose another post-hoc method that encodes each word in the text as a binary bit via a Bernoulli distribution ($p=0.5$), embedding the watermark through synonym substitution: words representing bit 0 are replaced with synonyms representing bit 1.
+Detection is again done via statistical testing.
+However, this method is fragile: synonyms are not reliably preserved under paraphrasing and often fail to capture subtle contextual meanings, which can noticeably degrade text quality and watermark robustness.
+Hao et al.
+(2025) is a post-hoc watermarking technique similar to Yang et al.
+(2023) that improves robustness by selecting semantically or syntactically essential words—those less likely to be altered during paraphrasing—as anchor points for embedding.
+The method uses paraphrase-based lexical substitution to insert watermarks while preserving the original semantics.
+However, empirical results in Chang et al.
+(2024) demonstrate that this method is not robust to paraphrasing compared to other methods such as KGW, SemStamp, and PostMark.
+Appendix B Experimental Settings
+In all combinations of the experiments, following Kirchenbauer et al.
+(2023), sampling from the LLM was performed with a temperature of 0.7 and a repetition penalty of 1.05, while the minimum and the maximum number of generated tokens were set to 195 and 205, respectively.
+The maximum number of rejection sampling iterations was set to 100, again to align with the code provided by Hou et al.
+(2024a, b).
+However, this setting reflects a trade-off between detection performance and generation speed.
+Based on our experiments, setting it to 25 achieves strong performance, with higher values offering only marginal improvements (see Appendix E).
+For token-level watermarking baselines, in cases where results were not directly extracted from Hou et al.
+(2024a, b), we employed the open-source MarkLLM watermarking framework Pan et al.
+(2024), with their recommended configurations ($\gamma=0.5,\delta=2$, prefix_length=1, etc.) to run the experiments.
+The majority of the experiments, including text generation and detection tasks, were conducted on a workstation equipped with an Intel Core i9 processor, 64GB of RAM, and an Nvidia RTX 3090 GPU with 24GB of VRAM.
+Some of the experiments involving bigram paraphrasing were performed on compute nodes with an Nvidia V100 GPU with 32GB of VRAM.
+Appendix C Averaged Performance Across Datasets
+Figure 5: Detection performance averaged across three datasets.
+Left: ROC-AUC ↑ across different paraphraser variants.
+Right: TP@FP=5% ↑ under the same settings.
+These radar plots provide a holistic comparison of all watermarking algorithms across paraphrasing conditions, averaged across datasets.
+SimMark demonstrates consistently strong robustness, closely matching or outperforming state-of-the-art baselines, particularly under heavier transformations.
+Dataset Algorithm No Paraphrase Pegasus Pegasus-Bigram Parrot Parrot-Bigram GPT3.5 GPT3.5-bigram Avg. Paraphrased
+Avg. Over Datasets
+UW 99.9 / 99.5 / 99.9 98.4 / 82.9 / 95.0 96.9 / 64.3 / 87.3 95.8 / 62.5 / 83.6 95.1 / 56.2 / 80.9 92.3 / 36.9 / 66.2 84.7 / 19.8 / 43.8 93.8 / 53.8 / 76.1
+KGW 99.5 / 98.3 / 98.7 95.8 / 86.3 / 91.2 93.4 / 55.5 / 75.3 87.5 / 32.1 / 58.2 86.3 / 26.7 / 53.6 84.8 / 20.6 / 49.9 77.3 / 9.9 / 31.6 87.5 / 38.5 / 60.0
+SIR 99.8 / 98.8 / 99.9 92.5 / 69.1 / 82.9 92.9 / 58.5 / 78.4 92.3 / 50.2 / 70.6 93.1 / 51.9 / 72.7 84.1 / 26.5 / 42.6 79.3 / 15.4 / 37.0 89.0 / 45.3 / 64.0
+SemStamp 99.5 / 96.6 / 98.0 98.4 / 90.3 / 94.8 97.7 / 85.4 / 91.7 93.9 / 65.1 / 79.2 95.1 / 70.6 / 82.9 86.9 / 42.3 / 57.8 85.3 / 37.5 / 52.8 92.9 / 65.2 / 76.5
+$k$-SemStamp 99.8 / 98.6 / 99.1 99.4 / 93.4 / 96.9 99.0 / 90.5 / 95.6 98.1 / 82.5 / 91.7 98.2 / 83.6 / 91.1 93.2 / 60.6 / 77.4 92.3 / 57.4 / 73.8 96.7 / 77.8 / 87.7
+Cosine-SimMark (ours) 99.5 / 97.3 / 98.6 99.2 / 92.7 / 97.7 99.1 / 92.7 / 97.5 98.8 / 89.4 / 97.1 98.9 / 88.7 / 97.1 96.9 / 70.2 / 89.4 94.3 / 51.1 / 78.8 97.8 / 80.8 / 92.9
+Euclidean-SimMark (ours) 99.9 / 99.1 / 99.5 98.3 / 83.2 / 93.9 98.2 / 80.8 / 92.3 97.1 / 70.5 / 87.7 97.1 / 69.6 / 87.8 97.5 / 76.6 / 90.1 94.8 / 62.5 / 79.5 97.2 / 73.9 / 88.6
+Table 3: Detailed average detection performance across the RealNews, BookSum, and Reddit-TIFU datasets under different paraphrasers.
+Each cell reports ROC-AUC ↑ / TP@1%FP ↑ / TP@5%FP ↑.
+Higher values indicate better performance across all metrics.
+Bold and underlined numbers denote the highest and second-highest values, respectively.
+SimMark consistently ranks among the top-performing methods in robustness across various paraphrasers averaged across datasets.
+For a comprehensive comparison across all paraphrasing settings (as shown in Table 1), including ROC-AUC and TP@FP thresholds (1% and 5%), we report the averaged detection performance metrics across datasets in Table 3.
+To better visualize relative performance trends across paraphrasers, we also present radar plots in Figure 5, showing aggregated TP@FP=5% and ROC-AUC scores (see Figure 3 in the main text for aggregated TP@FP=1% scores).
+These visualizations complement the table by providing an intuitive view of robustness across settings.
+Together, these summaries highlight that SimMark consistently outperforms or matches the baseline methods across diverse conditions.
+Appendix D Additional Experimental Results
+To demonstrate the model-agnostic nature of SimMark, we applied our algorithm to the Gemma3-4B model [^1] We employed google/gemma-3-4b (4B) model.
+Team et al.
+(2025).
+We evaluated both Cosine-SimMark and Euclidean-SimMark under different paraphrasing models across the same three datasets as before: RealNews subset of C4, BookSum, and Reddit-TIFU.
+The effectiveness and robustness of watermarking techniques can depend heavily on the characteristics of the underlying LLM and the nature of the generated text.
+To maintain consistency and reliability across experiments, we made the following modifications:
+  • Predefined Interval Adjustment: The sentences’ embedding similarity distribution under Gemma3-4B differed from those in OPT-1.3B, requiring new intervals.
+We set the predefined interval to [0.86, 0.90] for cosine similarity (without PCA), and [0.11, 0.16] for Euclidean distance (with PCA).
+  • Threshold Transferability: In contrast to our earlier experiments—where the detection threshold $\beta$ was determined per dataset—we fixed $\beta$ across all datasets in these experiments.
+Specifically, we determined the threshold using only non-watermarked data from the BookSum dataset, and then applied it across all three datasets without modification.
+This approach simulates a more realistic setting where the detector is calibrated on a single corpus but expected to generalize to others.
+The results demonstrate that our method maintains high detection performance even under this general configuration.
+  • Longer Generations: Since Gemma3-4B tends to generate longer sentences compared to OPT-1.3B model that we employed earlier, we increased the number of generated tokens from 200 to 300 to ensure a sufficient number of sentences for reliable hypothesis testing.
+Table 4 reports the detection performance in terms of ROC-AUC ↑ / TP@1%FP ↑ / TP@5%FP ↑, for each setting (↑: higher is better).
+Across all datasets and paraphrasing scenarios, SimMark remains highly effective, with both cosine similarity and Euclidean distance variants maintaining strong ROC-AUC and TP rates.
+These results affirm that SimMark maintains its performance across different LLM families (i.e., OPT and Gemma3) and datasets/domains, further validating the general applicability of our proposed algorithm.
+Dataset Method No Paraphrase Pegasus Pegasus-Bigram Parrot Parrot-Bigram Avg. Paraphrased
+RN
+Cosine-SimMark 99.5 / 96.2 / 97.9 93.5 / 57.9 / 75.6 93.0 / 54.9 / 73.9 92.3 / 50.1 / 71.7 92.1 / 49.2 / 72.8 92.7 / 53.0 / 73.5
+Euclidean-SimMark [^7] PCA is applied. 99.5 / 97.4 / 98.2 93.2 / 65.3 / 79.5 91.8 / 61.8 / 78.5 91.9 / 58.0 / 73.9 91.5 / 58.7 / 72.8 92.1 / 61.0 / 76.2
+BS
+Cosine-SimMark 99.9 / 99.6 / 99.8 97.7 / 75.8 / 90.5 97.4 / 73.3 / 90.0 97.0 / 67.8 / 87.6 96.9 / 67.7 / 86.7 97.2 / 71.1 / 88.7
+Euclidean-SimMark [^7] PCA is applied. 99.9 / 99.7 / 99.9 97.6 / 82.0 / 91.4 97.4 / 77.9 / 89.9 97.4 / 78.0 / 91.2 91.5 / 58.7 / 72.8 96.0 / 74.2 / 86.3
+TIFU
+Cosine-SimMark 99.8 / 99.1 / 99.5 97.4 / 78.8 / 91.5 97.0 / 76.6 / 89.3 89.4 / 32.8 / 61.7 90.7 / 36.6 / 63.8 93.6 / 56.2 / 76.6
+Euclidean-SimMark [^7] PCA is applied. 99.6 / 98.8 / 99.1 97.3 / 82.0 / 91.4 96.9 / 81.0 / 90.1 92.2 / 53.6 / 73.9 92.8 / 58.4 / 76.0 94.8 / 68.8 / 82.8
+Table 4: Performance of SimMark using Gemma3-4B model across paraphrasers and datasets (RealNews denoted as RN, BookSum denoted as BS, and Reddit-TIFU denoted as TIFU).
+Each cell reports AUC $\uparrow$ / TP@FP=1% $\uparrow$ / TP@FP=5% $\uparrow$ (↑: higher is better).
+The results demonstrate that SimMark maintains strong performance across different datasets and paraphrasing conditions, highlighting its robustness and model-agnostic nature.
+Appendix E Effect of Sampling Budget on Detection Performance
+To better understand the trade-off between generation speed and detection performance, we analyze the impact of the max_trials hyperparameter, which defines the upper limit on the number of rejection sampling iterations during watermark injection.
+While we set this value to 100 in our main experiments (to align with prior works of Hou et al.
+(2024a, b)), it is important to examine whether such a large value is necessary.
+Figure 6 shows two evaluation metrics—ROC-AUC ↑ and TP@FP=1% ↑ (↑: higher is better)—on the RealNews dataset for cosine-SimMark and OPT-1.3B model under different values of max_trials.
+As shown in the plots, performance improves significantly when increasing max_trials from 5 to 25, but plateaus thereafter.
+In particular, both ROC-AUC and TP@FP=1% show diminishing returns beyond 25 trials, indicating that additional sampling brings little performance gain.
+These results suggest that setting max_trials to 25 achieves a good balance between robustness and efficiency, and using larger values (e.g., 100) is not strictly necessary in practice.
+These findings, together with the average sampling statistics reported in the main paper (e.g., 7.1 samples per sentence), highlight SimMark’s ability to balance robustness with generation speed compared to SemStamp (20.9 samples per sentence) and $k$-SemStamp (13.3 samples per sentence).
+Figure 6: Impact of the maximum number of rejection sampling trials on detection performance.
+Increasing max_trials improves both ROC-AUC ↑ and TP@1%FP ↑ (↑: higher is better), but the improvement plateaus around 25.
+Results are reported on the RealNews dataset using cosine-SimMark and OPT-1.3B model.
+Appendix F Empirical Analysis of Generation Efficiency
+We experimentally evaluate the generation-time latency introduced by SimMark.
+Using 100 samples from the BookSum dataset, we measure the per-sentence generation time under various configurations of max_trials, using OPT-1.3B model.
+On our hardware (see Appendix B), with max_trials = 100 (i.e., maximum number of rejection sampling), cosine-SimMark yields an average latency of 1.33 seconds per sentence—corresponding to a 7.1× slowdown compared to the unwatermarked baseline average (0.188 sec/sentence).
+This matches our reported average sampling rate of 7.1 candidates per sentence (see Section 4).
+For reference, token-level methods such as UW and SIR yield 0.182 and 0.377 seconds per sentence, respectively (i.e., around 1× and 2× slowdown).
+This highlights the key trade-off between efficiency and robustness: while token-level methods are faster, SimMark offers less impact on text quality and stronger resistance to paraphrasing.
+Method Latency Slowdown
+Baseline (no watermark) 0.188 sec/sent 1.0×
+SimMark (max_trials=25) 0.977 sec/sent 5.2×
+SimMark (max_trials=50) 1.114 sec/sent 5.9×
+SimMark (max_trials=100) 1.330 sec/sent 7.1×
+UW (token-level) 0.182 sec/sent $\sim$1.0×
+SIR (token-level) 0.377 sec/sent 2.0×
+Table 5: Latency comparison of different watermarking methods (in seconds per sentence) and their slowdowns relative to unwatermarked generation.
+SimMark’s higher overhead stems from rejection sampling but remains practical—especially at lower max_trials—and can be further reduced with efficient decoding backends (e.g., vLLM).
+In table 5, we also evaluate SimMark under smaller max_trials values for practical trade-offs, as performance remains strong even with a smaller sampling budget (see Appendix E).
+While SimMark introduces overhead due to rejection sampling, newer decoding backends (e.g., vLLM Kwon et al.
+(2023)) can reduce this cost substantially via prompt-prefix reuse and KV cache optimization.
+However, our current benchmarks were conducted using the Hugging Face Transformers library Wolf et al.
+(2020) without these optimizations.
+Although we were unable to directly benchmark SemStamp and $k$-SemStamp due to technical issues, their substantially higher sampling requirements suggest that they would incur 2–3× greater overhead than SimMark.
+It is worth noting that a generation latency of about 1 second per sentence remains below the average human reading speed (200–250 wpm, or roughly 3–6 seconds per sentence).
+In summary, we believe SimMark offers a reasonable trade-off between detection robustness and runtime overhead, maintaining practicality for many real-world applications even in long-form scenarios.
+Appendix G Theoretical Analysis of Sampling Efficiency
+The average number of samples required to generate a valid sentence is influenced by the chosen interval $[a,b]$.
+To provide further insights into this relationship, we estimated the area under the curve (AUC) of the embedding similarity distribution for an unwatermarked LLM (OPT-1.3B).
+For instance, for the interval $[0.68,0.76]$ (for cosine-SimMark), the estimated AUC is approximately $0.194$.
+Using the mean of the geometric distribution, which is given by $\frac{1}{p}$, where $p$ is the probability of success (in this case, the probability of falling within the interval), this translates to an expected average of $\frac{1}{0.194}\approx 5.1$ samples per valid sentence.
+This estimate is on par with the experimental results reported in the main paper.
+The AUC estimates were computed using the binning technique, as described in Appendix L.
+This analysis underscores the importance of carefully selecting the interval $[a,b]$, as narrower intervals may increase the number of required samples, leading to reduced sampling efficiency but better performance, while broader intervals may compromise the effectiveness of the watermark.
+By understanding the interplay between the interval choice and sampling efficiency, we can better optimize SimMark’s performance.
+Appendix H Examples of Watermarked Text
+Figures 7 and 8 provide examples of text generated with and without the SimMark watermark using OPT-1.3B.
 These examples illustrate the imperceptibility of the watermark to human readers while enabling robust detection through our proposed algorithm.
 They also highlight SimMark’s robustness to paraphrasing while maintaining quality comparable to non-watermarked text.
-Figure 4: Example of text generated with and without cosine-SimMark using the RealNews dataset.
+Figure 7: Example of text generated with and without cosine-SimMark using RealNews dataset and OPT-1.3B model.
 The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid.
 Numbers in parentheses represent the soft count for partially valid sentences.
-The top panel shows non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=0.14\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@cmyk@stroke{0}{1}{1}{0}\pgfsys@color@cmyk@fill{0}{1}{1}{0}<}\;5.033$, false negative).
-The middle panel demonstrates text generated using SimMark with cosine similarity-based watermarking, producing a strong detection signal ($z_{\text{soft}}=9.48\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;5.033$).
-The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite semantic alterations ($z_{\text{soft}}=6.94\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;5.033$).
-Figure 5: Example of text generated with and without Euclidean-SimMark using the BookSum dataset.
+The top panel shows non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=0.14\;{\color[rgb]{0,0,0}<}\;5.033$, false negative).
+The middle panel demonstrates text generated using SimMark with cosine similarity-based watermarking, producing a strong detection signal ($z_{\text{soft}}=9.48\;{\color[rgb]{0,0.5,0}>}\;5.033$).
+The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite paraphrasing ($z_{\text{soft}}=6.94\;{\color[rgb]{0,0.5,0}>}\;5.033$).
+Figure 8: Example of text generated with and without Euclidean-SimMark using BookSum dataset and OPT-1.3B model.
 The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid.
 Numbers in parentheses represent the soft count for partially valid sentences.
-The top panel shows the non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=-1.07\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@cmyk@stroke{0}{1}{1}{0}\pgfsys@color@cmyk@fill{0}{1}{1}{0}<}\;4.13$, false negative).
-The middle panel demonstrates text generated using SimMark with Euclidean distance-based watermarking, producing a strong detection signal ($z_{\text{soft}}=13.07\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;4.13$).
-The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite semantic alterations ($z_{\text{soft}}=11.99\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;4.13$).
-Appendix C Ablation Study on Soft Count Smoothness Factor $K$
+The top panel shows the non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=-1.07\;{\color[rgb]{0,0,0}<}\;4.13$, false negative).
+The middle panel demonstrates text generated using SimMark with Euclidean distance-based watermarking, producing a strong detection signal ($z_{\text{soft}}=13.07\;{\color[rgb]{0,0.5,0}>}\;4.13$).
+The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite paraphrasing ($z_{\text{soft}}=11.99\;{\color[rgb]{0,0.5,0}>}\;4.13$).
+Appendix I Ablation Study on Soft Count Smoothness Factor $K$
 Count Method K Cosine-SimMark Paraphrased cosine-SimMark Euclidean-SimMark Paraphrased Euclidean-SimMark
@@
 Regular Count $\infty$ 99.7 / 97.2 / 99.1 99.1 / 88.7 / 97.6 99.8 / 98.5 / 99.7 97.0 / 70.0 / 88.2
-[…]
+Table 6: Ablation study on the smoothness factor $K$ in soft counting (Eq.
+(1)) using the RealNews dataset, with Pegasus as the paraphraser.
+Metrics reported include ROC-AUC $\uparrow$, TP@FP=1% $\uparrow$, and TP@FP=5% $\uparrow$, from left to right.
+The last row ($K=\infty$) corresponds to regular counting with a step function in the interval $[a,b]$.
+A smoothness factor of $K=250$ provides a good balance between performance before and after paraphrase attacks for both cosine-SimMark and Euclidean-SimMark.
+Notably, while soft counting slightly reduces performance in the absence of paraphrasing, it demonstrates enhanced robustness against paraphrasing, yielding an increase across all metrics for Pegasus paraphraser and potentially larger gains against more advanced paraphrasers.
 In this section, we analyze the impact of the smoothness factor $K$ on the performance of SimMark.
@@
 For example, TP@FP=1% improves by 1.6–2.3% for Pegasus-paraphrased text when $K=250$, and the improvement is likely to be even more significant for stronger paraphrasers.
-Appendix D Ablation Study on Impact of PCA
-Table 4 presents the results of an ablation study investigating the impact of applying PCA to reduce the dimensionality of sentence embeddings across the RealNews, BookSum, and Reddit-TIFU datasets.
+Appendix J Ablation Study on Impact of PCA
+Table 7 presents the results of an ablation study investigating the impact of applying PCA to reduce the dimensionality of sentence embeddings across RealNews, BookSum, and Reddit-TIFU datasets.
 Metrics include ROC-AUC, and TP at fixed FP rates (FP=1% and FP=5%).
@@
 For Euclidean distance-based SimMark, applying PCA generally improves robustness against paraphrasing attacks across most datasets, except for the BookSum dataset.
-This improvement likely arises because reducing dimensionality helps mitigate noise in the embeddings, especially after paraphrase attack.
+This improvement likely arises because reducing dimensionality helps mitigate noise in the embeddings, especially after the paraphrasing attack.
 On the other hand, for cosine similarity-based SimMark, applying PCA reduces performance across all datasets.
@@
 Dataset Configuration No paraphrase Pegasus
-RealNews Cosine-SimMark (No PCA) 99.7 / 96.9 / 98.8 99.2 / 90.3 / 98.2
+RealNews
+Cosine-SimMark (No PCA) 99.7 / 96.9 / 98.8 99.2 / 90.3 / 98.2
 Cosine-SimMark (PCA) 99.6 / 96.9 / 99.1 92.1 / 33.8 / 71.2
@@
 Euclidean-SimMark (PCA) 99.8 / 98.5 / 99.2 97.2 / 72.3 / 88.9
-BookSum Cosine-SimMark (No PCA) 99.8 / 98.8 / 99.5 99.5 / 93.3 / 98.5
+BookSum
+Cosine-SimMark (No PCA) 99.8 / 98.8 / 99.5 99.5 / 93.3 / 98.5
 Cosine-SimMark (PCA) 100 / 99.9 / 99.9 98.7 / 87.3 / 95.1
@@
 Euclidean-SimMark (PCA) 99.9 / 99.3 / 99.5 97.4 / 69.8 / 88.6
-Reddit-TIFU Cosine-SimMark (No PCA) 99.1 / 96.3 / 97.6 98.9 / 94.5 / 96.4
+Reddit-TIFU
+Cosine-SimMark (No PCA) 99.1 / 96.3 / 97.6 98.9 / 94.5 / 96.4
 Cosine-SimMark (PCA) 99.7 / 98.8 / 99.3 96.6 / 78.9 / 89.3
@@
 Euclidean-SimMark (PCA) 99.8 / 98.7 / 99.2 99.0 / 94.7 / 97.6
-[…]
-Appendix E Finding an Optimal Interval
+Table 7: Ablation study on the impact of applying PCA to embeddings across three datasets.
+Metrics reported include ROC-AUC $\uparrow$, TP@FP=1% $\uparrow$, and TP@FP=5% $\uparrow$, respectively, from left to right.
+Higher values indicate better performance across all metrics.
+Bold and underlined numbers denote the highest and second-highest values, respectively.
+For cosine-SimMark, not applying PCA yields better results, while for Euclidean-SimMark, applying PCA improves performance except on the BookSum dataset.
+Appendix K Finding an Optimal Interval
 Figure 9 shows the distribution of distances between embeddings of consecutive sentences for both human and LLM-generated text, calculated on a sample of size 1000 from the BookSum dataset (no PCA applied to the embeddings in this case).
 A small but noticeable distribution shift between the two can be observed.
-Based on this, the interval $[0.4,0.55]$ appears to be a reasonable choice for SimMark watermarking in this case.
+Based on this, the interval $[0.4,0.55]$ appears to be a reasonable choice for SimMark in this case.
 It is important to note that changes to the embedding representations, such as applying PCA or using a different embedding model, will lead to altered distance distributions.
@@
 Similarly, if we plot the figure for when cosine similarity is used instead of Euclidean distance, intervals $[0.81,0.94]$ and $[0.68,0.76]$ are good candidates for cases with and without PCA, respectively.
-This variability in the distance distributions also strengthens the algorithm’s resistance to reverse engineering.
+This variability in the distance distributions may also strengthen the algorithm’s resistance to reverse engineering.
 Selecting the optimal interval $[a,b]$ is a critical step in achieving a robust and reliable watermarking with SimMark.
@@
 Finally, further exploration of dynamic interval selection mechanisms could enhance SimMark’s robustness.
-Figure 6: Distribution of Euclidean distances between embeddings of consecutive sentences for human-written and LLM-generated text on the BookSum dataset, generated using OPT-1.3B.
+Figure 9: Distribution of Euclidean distances between embeddings of consecutive sentences for human-written and LLM-generated text on BookSum dataset, generated using OPT-1.3B.
 The figure demonstrates that the interval $[0.4,0.55]$ is a reasonable choice for Euclidean-SimMark in this case, though it is not necessarily the only viable option.
-Appendix F Computing Threshold $\beta$ for soft-$z$-test
+Appendix L Computing Threshold $\beta$ for soft-$z$-test
 Recall that a text is classified as LLM-generated when $z_{\mathrm{soft}}>\beta$, and as human-written otherwise.
@@
 This value of $p_{0}$ is a key component in calculating the $z_{\mathrm{soft}}$, as it represents the proportion of valid sentences in human-written text under the given interval.
-$p_{0}$ serves as an indicator of how frequently valid sentences are expected to occur in human-written text.
+$p_{0}$ serves as an indicator of how frequently valid sentences are expected to occur in human-authored text.
 To compute $p_{0}$, we analyze the distribution of similarities (or distances) using a histogram approach, such as the one depicted in Figure 9.
@@
 Mathematically, $p_{0}$ is estimated as:
-$\small p_{0}=\frac{\text{Number of samples in bins corresponding to }[a,b]}{\text{Size of the dataset}},$
+[…]
 once $p_{0}$ is estimated, the detection threshold $\beta$ is determined by iterating over a range of possible values, typically from -10 to 10, to find the one that results in the desired false positive rate.
 Specifically, the threshold is chosen such that the proportion of human-written texts misclassified as LLM-generated matches the target FP rate (e.g., 1% or 5%).
-It is worth noting that the distribution of similarities or distances may vary depending on different factors such as embedding model, and similarity measure (e.g., cosine or Euclidean).
+It is worth noting that the distribution of similarities or distances may vary depending on different factors such as the embedding model and similarity measure (e.g., cosine or Euclidean).
 As a result, $p_{0}$ and therefore $\beta$ are determined programmatically during the detection to ensure reliable performance of the watermarking algorithm.
-Appendix G Analysis of Sampling Efficiency
-The average number of samples required to generate a valid sentence is influenced by the chosen interval $[a,b]$.
-To provide further insights into this relationship, we estimated the area under the curve (AUC) of the embedding similarity distribution for an unwatermarked LLM.
-Specifically, for the interval $[0.68,0.76]$ (for cosine-SimMark), the estimated AUC is approximately $0.194$.
-Using the mean of the geometric distribution, which is given by $\frac{1}{p}$, where $p$ is the probability of success (in this case, the probability of falling within the interval), this translates to an expected average of $\frac{1}{0.194}\approx 5.1$ samples per valid sentence.
-This estimate is on par with the experimental results observed in our study.
-The AUC estimates were computed using the binning technique, as described in Appendix F.
-This analysis underscores the importance of carefully selecting the interval $[a,b]$, as narrower intervals may increase the number of required samples, leading to reduced sampling efficiency but better performance, while broader intervals may compromise the effectiveness of the watermark.
-By understanding the interplay between the interval choice and sampling efficiency, we can better optimize SimMark’s performance for different embedding models.
-Appendix H Experimental Environment
-The majority of the experiments, including text generation and detection tasks, were conducted on a workstation equipped with an Intel Core i9 processor, 64GB of RAM, and an Nvidia RTX 3090 GPU with 24GB of VRAM.
-Some of the experiments involving bigram paraphrasing were performed on compute nodes with Nvidia V100 GPU with 32GB of VRAM.
-Appendix I Prompts Used with GPT-3.5-Turbo
-Table 5 presents the prompts we used to obtain paraphrases using GPT-3.5-Turbo (accessed via OpenAI API [^8] https://platform.openai.com/docs/api-reference) for both regular paraphrasing and more aggressive bigram paraphrasing attacks.
-By using the same prompts as Hou et al.
-(2023), we ensured that our results were directly comparable to those extracted from their paper, maintaining consistency in evaluation methodology.
+Appendix M Human Evaluation of SimMark’s Imperceptibility
+Figure 10: Screenshot of the instructions and consent form shown to participants before the human evaluation study.
+It describes the task, evaluation criteria, expected time commitment, and privacy assurances.
+Human evaluation provides a more direct measure of imperceptibility for our watermarking algorithm.
+To assess this, we conducted a small-scale A/B test comparing outputs generated by SimMark to unwatermarked ones, using 10 randomly selected BookSum samples.
+Each sample pair consisted of one watermarked and one unwatermarked text.
+Option Chosen Percentage
+SimMark identified as watermarked 36.7%
+No noticeable difference 23.3%
+Unwatermarked identified as watermarked 40.0%
+Table 8: Results of a small-scale human evaluation measuring the imperceptibility of SimMark.
+Participants were asked to identify which of two outputs (watermarked vs.
+unwatermarked) appeared to be watermarked across 10 random samples.
+Responses were nearly evenly split, indicating that SimMark watermarking is largely imperceptible to human readers.
+Three volunteers (all master’s students in Canada, Iranian, aged between 18 and 25, with a Computer Science background) participated in the study.
+No sensitive data was collected beyond age range, academic level, and general field of study.
+Before starting the evaluation, participants were shown the consent and instruction form (Figure 10), which described the task, the absence of any risks, and the use of their responses solely for academic purposes.
+Participants provided explicit consent by filling out the form and then proceeded with the evaluation task.
+After filling in the form, participants were asked to identify which option—A or B—was more likely to contain a watermark.
+We also included the “N” (no noticeable difference) option to avoid forcing a binary choice when participants genuinely could not distinguish between the texts, thereby making the evaluation more reliable given the scale of the evaluation.
 Prompt for Regular Attack
-[…]
+Previous ${context} \n$ Current sentence to paraphrase: {sent}
 Prompt for Bigram Attack
-[…]
+Previous ${context} \n$ Paraphrase in {num_beams} different ways and return a numbered list: {sent}
 Table 9: Prompts used to generate paraphrases with GPT-3.5-Turbo for regular and bigram attacks.
 These are the same prompts used by Hou et al.
-(2023) for consistent and comparable evaluation.
-Here, sent represents the target sentence to rephrase, context includes all preceding sentences, and num_beams specifies the number of paraphrases generated for the bigram attack.
-A higher num_beams value indicates a more aggressive attack.
+(2024a) for consistent and comparable evaluation.
+Here, “sent” represents the target sentence to rephrase, “context” includes all preceding sentences, and “num_beams” specifies the number of paraphrases generated for the bigram attack.
+A higher “num_beams” value indicates a more aggressive attack.
+Following Hou et al.
+(2024a), we set “num_beams” to 10 to have 10 rephrases of each sentence.
+We focused our human evaluation on asking participants to detect watermarking, rather than assess fluency or preference.
+This directly targets SimMark’s design objective of imperceptibility: ensuring that even when readers are explicitly prompted to spot a watermark, they find it difficult to do so.
+As shown in Table 8, on average, participants labeled the SimMark output as watermarked in 36.7% of cases, while in 40.0% of cases they mistakenly identified the unwatermarked text as watermarked.
+In 23.3% of comparisons, they reported no noticeable difference.
+These results are close to random guess, suggesting that SimMark’s watermark is largely imperceptible to human readers while maintaining natural fluency.
+A larger-scale evaluation is deferred to future work.
+Appendix N Prompts Used with GPT-3.5-Turbo for Paraphrasing
+Table 9 presents the prompts we used to obtain paraphrases using GPT-3.5-Turbo (accessed via OpenAI API [^2] https://platform.openai.com/docs/api-reference) for both regular paraphrasing and more aggressive bigram paraphrasing attacks [^3] Used “gpt-3.5-turbo-16k” model..
+By using the same prompts as Hou et al.
+(2024a), we ensured that our results were directly comparable to those extracted from their paper, maintaining consistency in evaluation methodology.